### v2

# Demystifying Misconceptions in Social Bots Research

Demystifying Misconceptions in Social Bots Research
Stefano Cresci,1,∗ Kai-Cheng Yang,2 Angelo Spognardi,3 Roberto Di Pietro,4
Filippo Menczer,5 Marinella Petrocchi1
1IIT-CNR, Italy
2Northeastern University, USA
3“La Sapienza” University of Rome, Italy
4KAUST, Saudi Arabia
5Indiana University, USA
Research on social bots aims at advancing knowledge and providing solutions to one of the most debated forms of online manipulation. Yet, social bot research is plagued by widespread biases, hyped results, and misconceptions that set the stage for ambiguities, unrealistic expectations, and seemingly irreconcilable findings. Overcoming such issues is instrumental towards ensuring reliable solutions and reaffirming the validity of the scientific method. In this contribution, we review some recent results in social bots research, highlighting and revising factual errors as well as methodological and conceptual biases. More importantly, we demystify common misconceptions, addressing fundamental points on how social bots research is discussed. Our analysis surfaces the need to discuss research about online disinformation and manipulation in a rigorous, unbiased, and responsible way. This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
Corresponding author: stefano.cresci@iit.cnr.it
Introduction
Human decision-making processes depend on the availability of high quality information. In addition, a healthy society requires a shared understanding of issues and values to foster trust and cooperation. Misinformation and disinformation erode trust and emphasize divisions. Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach consensus and take effective collective action. The repercussions could be so severe as to delay responses to a deadly pandemic (Cinelli et al., 2020; Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
The science of misinformation seeks solutions to these problems (Lazer et al., 2018). However, this paramount endeavor can itself incur in the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023). For example, scientific articles and publishers engage in a fierce competition for attention, similarly to what is being experienced by news outlets. As a consequence, sensationalist claims and hyped results can sometimes be used as shortcuts to publication and other scientific accomplishments (West and Bergstrom, 2021). Then, confirmation bias is making its way into science in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021). And again, over-generalizations of scientific results and poor understanding of methodological and conceptual limitations give rise to multiple misconceptions about misinformation (Altay et al., 2023). These and other problems currently undermine the efficacy and credibility of our research efforts. Therefore, for the science of misinformation to benefit our society, we must first solve the problems within the science itself.
The present study concerns one of the many forms of online disinformation: malicious social bots. Social bots are accounts on an online platform that are—at least partially—automated (Ferrara et al., 2016; Cresci, 2020). Thanks to their ease of use and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content. Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb misinformation and disinformation (Shao et al., 2018). For this reason, social bots have attracted considerable scholarly and media attention (Allem and Ferrara, 2018; Assenmacher et al., 2020; González-Bailón and De Domenico, 2021; Chen et al., 2021). One of the latest such episodes was the public dispute on the real number of bots on Twitter that took place before Elon Musk’s acquisition of the platform (Varol, 2023). However, despite many years of research, the science of social bots is replete with the same problems that also plague the science of misinformation. These originate from uncorrected factual errors, from the presence of biases in how scientific results are cited and discussed, and from a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
In our present analyses, we focus on some recently published articles about social bots. A thorough analysis of their methods, assumptions, results, and conclusions reveals several pitfalls ranging from factual inaccuracies and methodological flaws to more subtle incidents such as misrepresentation of the existing literature and conceptual errors. Our discussion links these issues to common misconceptions in social bots and misinformation research. Our critical analysis thus serves a twofold goal. On the one hand, we highlight and revise factual errors, methodological biases, and conceptual issues. On the other hand, demystifying common misconceptions allows us to address fundamental issues on how science is produced and discussed (West and Bergstrom, 2021). Overcoming such issues is instrumental towards ensuring the credibility of science.
Methodological issues
Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Ferrara et al., 2016; Cresci, 2020). In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots—on various online platforms. The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts with the goal of detecting subtle differences between authentic users and automated entities. Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
Information leakage
Information leakage in machine learning refers to situations where a model is exposed to confidential information from the training data that is later exploited to make unfair inferences (Chakrabarti et al., 2008). Feature leakage is a specific type of information leakage where the model exploits machine learning features that are not available at prediction time. This issue typically results in the trained model exhibiting unrealistically good performance due to the exploitation of “forbidden” information. In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance. For example, each botnet typically exhibits some peculiar characteristics that surface as a result of how the bots were created, their goals, or any other shared characteristic between them (Zhang et al., 2016; Mazza et al., 2022). These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Barabasi, 2005; Cresci et al., 2020). However, most of the time, such peculiar characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets. Therefore, if a bot detector exploits such characteristics at training time for classifying instances of known bots, it might learn spurious correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
Another way in which feature leakage occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise). Again, the correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading. Consequently, a detector exploiting these features would achieve excellent performance on known data but would likely exhibit poor capacity to generalize to new data with different characteristics.
The inadvertent or deliberate inclusion of information leakage in a bot detector’s training process introduces a potential avenue for its authors to artificially elevate its performance, enabling strategic competition with established state-of-the-art models. Such situation can lead to the proposal of ostensibly superior models, which however represent a regression rather than advancement in addressing the bot detection problem. In addition, the same models could also be exploited to unduly undermine the performance of other models. This was recently done by Gallwitz and Kreil (2022) and Hays et al. (2023), who trained simple but unrealistic bot detectors using spurious signals, such as whether an account was verified by Twitter, to exaggerate the limitations of more general state-of-the-art models.
Cherry-picking
Another challenge in bot detection arises from the multitude of disparate detectors and benchmark datasets (Cresci, 2020). This creates an environment where authors introducing new detectors can selectively engage in favorable comparisons. The issue is exacerbated by the difficulty of delineating a clear state-of-the-art in bot detection, as the sheer volume of existing models makes it arduous to discern the most effective ones. The lingering uncertainty around the performance of even established bot detectors (Rauchfleisch and Kaiser, 2020) makes it possible to cherry-pick competitors and evaluation scenarios, allowing proponents of a novel detector to demonstrate its superiority only against a small subset of detectors and datasets, conveniently omitting those against which it may not perform as well.
To make matters worse, despite the existence of high-quality research published in prestigious venues (Shao et al., 2018; Stella et al., 2018; González-Bailón and De Domenico, 2021), the heightened interest in social bots has attracted a wide spectrum of contributions, including some lower-quality works. For example, Gallwitz and Kreil (2022) selected non-representative samples (prominent public accounts, such as members of the US Congress and Nobel Prize Laureates) to challenge the accuracy of a widely-used bot detection tool, inflating its false-positive rate. Such misleading claims underscore the fallacy of cherry-picking evaluation data. Since bot detection methods, whether based on machine learning or on human annotation, have inherent error margins, accuracy results can be manipulated by carefully choosing test examples. This less rigorous body of work, when added to the literature, creates a reference that, while possibly lacking in quality, further facilitates the selective cherry-picking of competitors.
The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that benefit from information leakage or that overfit to specific evaluation datasets. [^1] We refer to overfitting when a model depends too much on irrelevant features of the training data, with the result that it performs well on the training data but generalizes poorly to unseen data (Bramer, 2007). State-of-the-art models are designed for greater generality, rendering direct comparisons with these simpler yet contextually specialized models inherently unfair. In a recent example of this kind, overfitted classifiers with as few as four features are compared to more complex and general state-of-the-art models (Hays et al., 2023).
The issue of cherry-picking extends beyond the unfair comparisons between bot detectors. A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings. Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023) and to falsely claim that bot detectors have not been systematically evaluated (Gallwitz and Kreil, 2022). These claims can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020). On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—to give a false impression of prior research supporting their claims. Such selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art in bot detection. By self-citing unvetted claims, authors create an illusion of authority. And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed to the same or comparable insights. This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of researchers who have previously advanced the field.
Addressing the cherry-picking issue necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature. Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts. This discrepancy surfaces a crucial tension in maintaining the quality and rigor of peer review processes within rapidly evolving research landscapes (Van Noorden, 2023). The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of questionable quality.
Straw-man methodology
The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation as if it faithfully represented the original research. For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on misrepresentation of the literature. As part of their criticism of supervised methods, they falsely claim that other researchers do not share or publish training and evaluation datasets. They also make demonstrably false claims about other researchers manipulating results through the use of unreported training data or threshold hacking. Gallwitz and Kreil criticize supervised bot detectors as invalid on the basis of these serious claims of scientific misconduct, yet do not provide any evidence in support of their allegation. Another methodological straw-man fallacy is the claim that Gallwitz and Kreil failed to reproduce an experiment by Keller and Klinger (2019). Gallwitz and Kreil applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in the bot detector.
A different manifestation of the straw-man fallacy is the claim that bot detection is exclusively a supervised machine learning task. On the one hand, this assumption could be motivated by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of available supervised bot detectors (Cresci, 2020). On the other hand, such a view overlooks a similarly robust literature highlighting that supervised detectors suffer from lack of generalizability and transferability, as well as other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). A variety of unsupervised approaches to bot detection have recently emerged as possible solutions to mitigate these problems (Cresci, 2020). Evaluating the generalization capabilities of bot detectors is precisely the goal of Hays et al. (2023). However, they chose to do so by adopting a supervised learning methodology that has been proven to suffer from critical issues when it comes to generalizability and transferability. Another issue with their methodology is that it does not fully support their conclusions. Specifically, Hays et al. leverage their results to criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods. They do not make any distinction between the different types of bot detectors proposed to date—supervised ones based on the analysis of individual accounts vs. unsupervised ones based on the analysis of groups of accounts (Cresci, 2020). This is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable. However, the analysis only evaluates supervised detectors, which have known generalization issues (Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021). The unsurprising—and biased—result is that the considered detectors fail to generalize, which is used to criticize the whole field.
Data biases
In machine learning and data science tasks, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009). As mentioned above, numerous datasets of bot and human accounts have been published over the years. On the one hand, this made it easier to thoroughly test the performance of new bot detectors. On the other hand, it introduced the risk of cherry-picking favorable datasets, potentially distorting reported performances and introducing bias. Furthermore, the temporal heterogeneity stemming from the availability of datasets spanning varying time periods—ranging from recent to decade-old—can pose additional challenges when detectors are trained on outdated data, undermining the relevance of their performance in contemporary settings. When considering the representativeness of published benchmark datasets with respect to the bots that currently inhabit online platforms, the readers should be aware that the accounts included in a dataset are likely a few years older than the publication date of the dataset itself. Moreover, some datasets are built by combining or extending previously published resources. For example, the pan-2019 dataset (Rangel and Rosso, 2019) is largely based on other datasets published between 2011 and 2017. Hence, the publication date of a dataset should be considered as a generous upper bound of the recency of the accounts therein. Even under this relaxed assumption, in consideration of the rapid pace of change of the landscape of online harms and of the availability of more comprehensive and more recent datasets (Feng et al., 2021), we may run the risk of capturing datasets that are no longer representative of the actual state of the platforms. This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021).
Some recent literature investigates the characteristics of accounts that are over ten years old. Obsolete datasets can be used to criticize other people’s detection methods or to promote one’s own method as better than the state of the art. A glaring example is the work of Hays et al. (2023), where out of eleven analyzed datasets, the most recent was published in 2020 and the least recent in 2011.
Annotations can be another source of data bias. For example, so-called cyborg accounts (Ferrara et al., 2016) are automated in part via software, but closely managed and monitored by humans who can step in as needed to avoid detection. By labeling such accounts as human on the basis that they can reply to a query and assert that they are not bots, one can make it appear that a bot detection method has a higher false-positive rate than it actually does. To avoid this kind of bias, account labeling should be performed by multiple annotators following a shared and published rubric. Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) that accounts exhibit should be favored in place of binary labels.
Some work in the literature fails to follow this methodology, only to accuse other researchers of bias. Gallwitz and Kreil (2022) demonstrate this behavior in several cases. Without disclosing any annotation rubric, they change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative. In one case, criticizing Yang et al. (2019), they label accounts that automatically post news headlines, such as those associated with major newswire agencies, as human. Their criticism of Keller and Klinger (2019) goes further: first, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger; second, they manually inspected only 0.04% of this different set; third, they ignored suspended accounts; and, fourth, they labeled as human accounts those that cross-posted tweets through software apps. Their conclusion was that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word.» Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al. (2020): suspended accounts are again assumed to be human to boost the alleged false-positive rate, then accounts that are clearly automated are still labeled as classification errors by changing the standard from “bot” to “malicious” or “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions is a dangerous trend toward scientific manipulation.
(a) Simplistic and naïve model of social bots and human-operated accounts. Here, all bots are alike, they are evenly distributed, and we have complete knowledge on their numbers and characteristics. According to this model, collecting unbiased and comprehensive bot datasets is possible.
(b) Realistic model of social bots and human-operated accounts, and of the related knowledge gaps. Here, social bots are organized in botnets, each with peculiar characteristics (color-coded). Among the missing information is the number of bots in known botnets, as well as the number, size, and characteristics of all unknown botnets.
Figure 1: The ecology of online accounts: social bots and humans. Differences between naïve and realistic models, and their implications for data collection.
Conceptual issues
Beyond methodological challenges, the landscape of social bots research is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues. These nuanced challenges, if left unaddressed, have the potential to exert an even more insidious influence on the development of the field and its reception.
Failure to account for context
In discussing the limitations that currently hinder progress in bot detection, some recent research has pointed the finger at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023). This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in datasets acquired with rigorous practices, contrary to the literature so far. More specifically, this criticism of the existing data collection practices subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient. Unbiased datasets are needed so as to avoid that peculiar bot and human characteristics leak into the data allowing simple models to achieve good performance on the collected data, while showing poor performance in real scenarios. Furthermore, comprehensive datasets are needed so that expressive models can learn to generalize to all types of existing social bots. Unfortunately, the assumption of the existence of an unbiased and comprehensive dataset is fallacious.
Regarding bias, many bot datasets were found to contain biases, in that the accounts therein featured some peculiar characteristics (Hays et al., 2023). We note that this is largely unsurprising since such peculiar characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a single entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016). Given that all bots belonging to the same botnet are created and/or operated by the same entity, and pursue the same common goal, they tend to share some similarities. When compared to other accounts, the botnet thus appears to have some peculiar characteristics, as highlighted in Figure 1(b). Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather should be traced back to the very nature of the social bot phenomenon.
The assumption about the possibility to build comprehensive bot datasets is equally flawed. Creating such datasets would require a uniform and random sample of an adequate number of accounts from the complete distribution of existing bots. This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space. While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance. Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020). we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible. In addition to the many challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021; Torres-Lugo et al., 2022). Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023). We only have a partial understanding and visibility of the botnets that operate on our online platforms. [^3] The court dispute between Elon Musk and former Twitter CEO Parag Agrawal on the real number of bots on Twitter is a paramount example of the intrinsic difficulties and ambiguity of tracking social bots. Figure 1 provides a conceptual view of the landscape of social bots and human-operated accounts: firstly according to a naïve interpretation (Figure 1(a)) and then in a more realistic representation (Figure 1(b)). Figure 1(b) highlights the knowledge gaps related to the presence of bots in online platforms. These include known unknowns: the exact number of bots belonging to known botnets. But they also include unknown unknowns: information about the hidden botnets operating on a platform. How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms? This fundamental question cannot currently be answered.
Common misconceptions in social bots research
Part of the social bot literature claims that the bot detection task has been effectively solved. Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications. This optimistic view positions bot detection research as a success story for machine learning, highlighting the progress made in achieving practical and deployable solutions in the field.
We disagree with this interpretation of the state of social bot research. On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection. We are well aware of this body of work, having actively contributed to its development for many years. Below we report on some notable examples. The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019). The limitations of benchmark datasets for bot detection were discussed by Feng et al. (2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al. (2022). Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021). Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021). Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media. [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022). For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020; Pacheco et al., 2021). This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works. This discrepancy introduces a common misconception about social bots research:
Misconception 1 Social bot detection is a solved task.
The above evidence shows that despite significant efforts devoted for a prolonged time (Cresci, 2020), bot detection is nowhere near to being a solved problem—quite the contrary.
In framing the bot detection task as a seemingly straightforward problem, some authors also hint at the possibility of easily adapting existing detectors to overcome their current limitations and keep pace with the evolution of more human-like bots (Hays et al., 2023). This example introduces another misconception about bot detection:
Misconception 2 Bot detection performance can be improved easily.
For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Cresci et al., 2021). Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term. It is reasonable to assume that future advances will suffer a similar fate. Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology. Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
Below we discuss some other misconceptions that undermine our understanding of social bots and of the related literature:
Misconception 3 All social bots are similar.
Misconception 4 Each bot detector can detect all types of bots.
A common assumption in the scientific literature, as well as among the general public, is that if a given bot detector has achieved good detection performance against some bots, then that detector should be able to replicate that performance on any bot detection task. This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022) and only considering the known botnets. Consider, for example, the differences between the bots used to boost the popularity of certain public figures—so-called fake followers (Cresci et al., 2015)—and those that manipulate trending topics—astroturfers and spammers (Abokhodair et al., 2015). Or the differences between bots involved in political manipulation (Shao et al., 2018; Caldarelli et al., 2020) versus those aimed at fooling automated trading systems (Tardelli et al., 2022; Bello et al., 2023). In other words, not all bots are the same. Some scholars addressed this heterogeneity by designing detectors that aim for generality and broad applicability, such as Botometer, especially in its latest releases (Sayyadiharikandeh et al., 2020; Yang et al., 2020). Others have developed detectors specifically designed to detect certain types of bots. The latter trade generalizability and portability for detection accuracy. Their performance depends heavily on the characteristics of the bots they are designed to detect. For example, a detector designed to detect time-synchronized retweeting actions (Mazza et al., 2019) would likely be useless at detecting mass-following bots (Cresci et al., 2015). However, this should not be regarded as a limitation of the bot detector—let alone one deliberately unstated by its developers in order to pass it off as a good product—but rather as an inappropriate use of the detector itself. Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized detectors. In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot or not. Therefore, in general, any detector, specialized or otherwise, has variable performance and its detection capabilities depend on the characteristics of the accounts it is applied to. In conclusion, no single bot detector is capable of detecting all types of bots.
Many articles proposing a new bot detector, or investigating bot interference, motivate the work on the basis of the role of social bots in spreading disinformation (Shao et al., 2018). Similar arguments are made for the role of bots in exacerbating social conflicts (Stella et al., 2018). In other words, part of the social bot literature is based on the assumption that bots are responsible for many of the problems that affect our online platforms and the information ecosystem as a whole. This brings to the fore another misconception:
Misconception 5 Social bots are mainly responsible for the spread of disinformation.
An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots. For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024). The existing literature has focused almost exclusively on detecting bots and characterizing their behavior, leaving the fundamental task of measuring the impact of bot malfeasance largely unexplored (Cresci, 2020). For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users. What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020; Pierri et al., 2023). Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022). Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation. It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
The previous misconceptions might induce readers to think that social bot research has led to flawed—if not outright useless—results. Such a claim has been made by Hays et al. (2023). Similarly, Gallwitz and Kreil (2022) not only question the validity of existing bot detection methods—based on the fallacies discussed earlier—but go even further in arguing that social bots are not an issue. These claims lead to our last, but by no means least important, misconception:
Misconception 6 All social bots research is flawed and bot detection results are useless.
We see at least two strong arguments against this thesis. First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns. For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014). Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023). In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the findings. These cases are just some of the success stories of social bot research. If researchers had not developed bot detection techniques, we would not have been able to identify pockets of anomalous accounts engaged in sophisticated manipulations. Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none. Second, the benefits of social bot research extend beyond the detection of malicious bots. For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024). In addition, social bot research has contributed to the early development of other neighboring fields. The early research on social bots, which dates back to 2010 (Ratkiewicz et al., 2011; Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years. In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation. In light of these considerations, research on social bots—imperfect as it is—seems far from useless. In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
Challenges in the post-API era
In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies. Previously, studies primarily concentrated on Twitter/X, likely because its free API facilitated data acquisition. However, this accessibility ceased in 2023 when Twitter/X terminated free data access for researchers. [^7] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram. [^8] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence in real time. It significantly impedes the collection of new bot samples necessary for studying their characteristics and training novel machine learning classifiers. Even if researchers can develop new classifiers, deploying them at scale is challenging, exposing social media users to potential manipulation. Bot operators, on the other hand, remain largely unaffected. Leveraging burner and virtual cellphones, they can manage bot accounts across platforms directly, bypassing API restrictions. The advent of AI-powered social bots, which evade current bot detection models (Yang and Menczer, 2023; Yang et al., 2024), further emphasizes the importance of data availability in support of social bot research.
Despite these setbacks, Europe’s Digital Service Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests. [^9] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs. However, these programs are significantly limited compared to previous access levels. Their opaque and stringent application review processes also leave their efficacy in question. Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics. Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors. The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community.
Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
A call for moral responsibility
Our analysis of the drawbacks and limitations of some recent studies brought to light the need for responsibility when discussing results about social bots. The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion. To this end, it is paramount to avoid propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars—such as those between proponents and opponents of social bots research.
The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society. As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses. Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019). Indeed, we believe that all the literature that propagates the misconceptions demystified in this paper provides misleading contexts in some way. Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021). Similarly, while making hyped or sensationalist claims can help publish a paper and accrue citations, such claims also contribute to creating misleading contexts and unrealistic expectations, as well as to increasing scientific tribalism.
Making tangible progress in the science of misinformation requires that we address many conceptual, practical, and ethical challenges. Doing so involves embracing the intrinsic complexity of the phenomenon, considering multiple viewpoints, and providing nuance rather than naivety. Sweeping statements such as “the creators of bot datasets are responsible for the failure of the field,” “all social bots research is flawed or useless,” or even “this new bot detector has flawless performance” do not contribute to achieving these goals.
In conclusion, should we fail to respect our moral obligation, we would produce biased and unreliable research, further worsening the problems that currently undermine the credibility of our field (West and Bergstrom, 2021; Altay et al., 2023). In a 2019 piece published in Science, Derek Ruths compared contrasting findings about social bots, commenting that «research on misinformation has come to resemble the very thing it studies» (Ruths, 2019). It is our moral responsibility to reverse this fatal course. This can only be achieved via responsible research that fosters nuanced, unbiased, and balanced viewpoints, and by a review process that adheres to the same principles. This article aims to make a contribution in this direction by debunking common fallacious arguments adopted by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
Acknowledgements
This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; by the European Union – Horizon 2020 Program under the scheme “INFRAIA-01-2018-2019 – Integrating Activities for Advanced Communities”, Grant Agreement n. 871042, “SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics”; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.

### v3

@@
 Corresponding author: stefano.cresci@iit.cnr.it
-Introduction
-Human decision-making processes depend on the availability of high quality information.
+1 Introduction
+Human decision-making processes depend on the availability of high-quality information.
 In addition, a healthy society requires a shared understanding of issues and values to foster trust and cooperation.
 Misinformation and disinformation erode trust and emphasize divisions.
-Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach consensus and take effective collective action.
-The repercussions could be so severe as to delay responses to a deadly pandemic (Cinelli et al., 2020; Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
+Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach a consensus and take effective collective action.
+The repercussions could be so severe as to delay responses to a deadly pandemic (Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
 The science of misinformation seeks solutions to these problems (Lazer et al., 2018).
-However, this paramount endeavor can itself incur in the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023).
-For example, scientific articles and publishers engage in a fierce competition for attention, similarly to what is being experienced by news outlets.
+However, this paramount endeavor can itself incur the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023).
+For example, scientific articles and publishers engage in a fierce competition for attention, similar to what is being experienced by news outlets.
 As a consequence, sensationalist claims and hyped results can sometimes be used as shortcuts to publication and other scientific accomplishments (West and Bergstrom, 2021).
@@
 Therefore, for the science of misinformation to benefit our society, we must first solve the problems within the science itself.
-The present study concerns one of the many forms of online disinformation: malicious social bots.
-Social bots are accounts on an online platform that are—at least partially—automated (Ferrara et al., 2016; Cresci, 2020).
-Thanks to their ease of use and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content.
+The present study concerns one of the many forms of online manipulation: social bots.
+In this context, let us make a brief clarification of what, for the purposes of this article, is a bot, a social bot, and a malicious social bot.
+In literature, the term bot is associated with several others that are often, and sometimes erroneously, used as synonyms.
+As discussed by Gorwa and Guilbeault (2020b), scholars, journalists, and policymakers often speak indiscriminately of bots, chatbots, botnets, sybils, and cyborgs, to cite a few.
+If we go back in time, the bots better known as chatbots referred to early programs that can converse with humans.
+Later in time, the word bot also came to mean a computer machine compromised, co-opted, and remotely controlled by malware (Yang et al., 2014).
+When these types of accounts work together, they form a network called botnet, whose methods of operation and proposals for detection have been studied by information security researchers (Yang et al., 2014; Antonakakis et al., 2017).
+With the advent of social media and social networks, bots were also framed as compromised network accounts, called sybils in the information security field.
+Considering the bots that populate online social networks—conventionally called social bots (Ferrara et al., 2016)—there are works that categorize such accounts according their intent or their capacity to imitate humans (Stieglitz et al., 2018), as in the case of social bots promoting or suppressing certain political narratives (Pratelli et al., 2024) or facilitating the spread of disinformation (Shao et al., 2018).
+Others differentiated social bots according to their activity on the platform, identifying those that systematically retweet the posts of political candidates (Cresci et al., 2017), or those that mass-follow a brand or character to boost their popularity (Cresci et al., 2015).
+There even exist social bots with no malicious intent that rather serve a useful purpose for those who follow them (Cresci, 2020; Savage et al., 2016).
+For example, these bots can act as news aggregators (Lokot and Diakopoulos, 2016).
+Among the many diverse existing definitions, here we define bots as social media accounts that are totally or partially automated.
+By focusing on the use of automation, we adopt a definition that is both widely used and based on an objective criterion (Ferrara et al., 2016; Wischnewski et al., 2021; Ferrara, 2022).
+This is in contrast to those based on the intent (malicious, neutral, or benign), on the capacity to imitate humans, or on the type of activity carried out, which are inherently more subjective.
+Nonetheless, we clarify that independently of the chosen definition, the issues highlighted in the remainder of this article as well as the proposed solutions, are broadly valid and applicable, ensuring that adhering to a different definition does not diminish the overall validity of the analyses and discussions herein.
+Thanks to their automated nature, ease of use, and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content.
 Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb misinformation and disinformation (Shao et al., 2018).
@@
 Overcoming such issues is instrumental towards ensuring the credibility of science.
-Methodological issues
-Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Ferrara et al., 2016; Cresci, 2020).
+2 Methodological issues
+Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Cresci, 2020).
 In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots—on various online platforms.
@@
 Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
-Information leakage
-Information leakage in machine learning refers to situations where a model is exposed to confidential information from the training data that is later exploited to make unfair inferences (Chakrabarti et al., 2008).
-Feature leakage is a specific type of information leakage where the model exploits machine learning features that are not available at prediction time.
-This issue typically results in the trained model exhibiting unrealistically good performance due to the exploitation of “forbidden” information.
+Model drift
+Model drift happens when a model relies on a set of features that sensibly diverge between train and test samples.
+Model drift generally happens when data slowly change distribution over time.
+However, it can also happen when the training and prediction datasets are significantly different.
+Generally, model drift can slowly reduce model accuracy over time, but it can also misleadingly show that the trained model performs well due to the exploitation of feature peculiarities not found in the prediction dataset.
 In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance.
 For example, each botnet typically exhibits some peculiar characteristics that surface as a result of how the bots were created, their goals, or any other shared characteristic between them (Zhang et al., 2016; Mazza et al., 2022).
-These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Barabasi, 2005; Cresci et al., 2020).
+These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Cresci et al., 2020).
 However, most of the time, such peculiar characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets.
-Therefore, if a bot detector exploits such characteristics at training time for classifying instances of known bots, it might learn spurious correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
-Another way in which feature leakage occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise).
-Again, the correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading.
+Here it is important to recall that the end goal of any bot detector—as with any machine learning classifier—is to accurately detect unknown bots from a set of never-seen-before data.
+Thus, the detection performance on a held-out portion of known data serves merely as a proxy for the real-world detection performance in-the-wild.
+However, if a bot detector exploits known characteristics at training time for classifying instances of known bots, it might learn correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
+Another way in which a model drift occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise).
+The correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading.
 Consequently, a detector exploiting these features would achieve excellent performance on known data but would likely exhibit poor capacity to generalize to new data with different characteristics.
-The inadvertent or deliberate inclusion of information leakage in a bot detector’s training process introduces a potential avenue for its authors to artificially elevate its performance, enabling strategic competition with established state-of-the-art models.
 Such situation can lead to the proposal of ostensibly superior models, which however represent a regression rather than advancement in addressing the bot detection problem.
@@
 This was recently done by Gallwitz and Kreil (2022) and Hays et al.
-(2023), who trained simple but unrealistic bot detectors using spurious signals, such as whether an account was verified by Twitter, to exaggerate the limitations of more general state-of-the-art models.
-Cherry-picking
-Another challenge in bot detection arises from the multitude of disparate detectors and benchmark datasets (Cresci, 2020).
-This creates an environment where authors introducing new detectors can selectively engage in favorable comparisons.
+(2023), who trained simple but unrealistic bot detectors using features tightly bound to the training dataset, such as whether an account was verified by Twitter—which was among the features used to select and label the data—to exaggerate the limitations of more general state-of-the-art models.
+The use of a feature used to select and label the data can also be framed as a feature leakage issue: it is crucial to refrain from using a feature that is utilized to select and label the data as a training input for the model since it will eventually lead to unreliable and inaccurate predictions.
+Unfair evaluation
+Another challenge in bot detection arises from the multitude of disparate definitions, detectors, and benchmark datasets (Cresci, 2020).
+For example, a critical issue in evaluating bot detectors is the possible inconsistency between the bot definition assumed by the detector and that used by the evaluator.
+This inconsistency can lead to biased and unfair evaluations as it is conceptually flawed to evaluate a detector using a different definition than that employed by the detector itself.
+A tool trained to identify a specific type of bot should be assessed based on its ability to detect that specific type, not alternative ones for which it was not designed.
+This misalignment in definitions can result in evaluations that misrepresent the detector’s true performance and capabilities.
+Therefore, while it is perfectly acceptable to discuss—and even criticize—the social bot definition used by a detector, evaluators must ensure that the definition used to evaluate a detector aligns with the detector’s, or be very clear about the differences and their implications towards the results of the evaluation.
+We point out, however, that there exists a trade-off between the need for consistency in using the same definition as the detector and the necessity of assessing the detector’s generalizability, which requires testing it against bots with different characteristics (Cresci, 2020).
+Striking this balance depends on multiple factors, including the practical context in which the detector is used and whether the detector is specialized towards specific types of bots or general-purpose.
+In addition, the existence of many bot detectors creates an environment where authors introducing new detectors can selectively engage in favorable comparisons.
 The issue is exacerbated by the difficulty of delineating a clear state-of-the-art in bot detection, as the sheer volume of existing models makes it arduous to discern the most effective ones.
 The lingering uncertainty around the performance of even established bot detectors (Rauchfleisch and Kaiser, 2020) makes it possible to cherry-pick competitors and evaluation scenarios, allowing proponents of a novel detector to demonstrate its superiority only against a small subset of detectors and datasets, conveniently omitting those against which it may not perform as well.
-To make matters worse, despite the existence of high-quality research published in prestigious venues (Shao et al., 2018; Stella et al., 2018; González-Bailón and De Domenico, 2021), the heightened interest in social bots has attracted a wide spectrum of contributions, including some lower-quality works.
-For example, Gallwitz and Kreil (2022) selected non-representative samples (prominent public accounts, such as members of the US Congress and Nobel Prize Laureates) to challenge the accuracy of a widely-used bot detection tool, inflating its false-positive rate.
-Such misleading claims underscore the fallacy of cherry-picking evaluation data.
-Since bot detection methods, whether based on machine learning or on human annotation, have inherent error margins, accuracy results can be manipulated by carefully choosing test examples.
-This less rigorous body of work, when added to the literature, creates a reference that, while possibly lacking in quality, further facilitates the selective cherry-picking of competitors.
-The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that benefit from information leakage or that overfit to specific evaluation datasets.
+The evaluation landscape is further complicated by the seemingly excellent performance achieved by models that overfit to specific evaluation datasets.
 [^1] We refer to overfitting when a model depends too much on irrelevant features of the training data, with the result that it performs well on the training data but generalizes poorly to unseen data (Bramer, 2007).
-State-of-the-art models are designed for greater generality, rendering direct comparisons with these simpler yet contextually specialized models inherently unfair.
-In a recent example of this kind, overfitted classifiers with as few as four features are compared to more complex and general state-of-the-art models (Hays et al., 2023).
-The issue of cherry-picking extends beyond the unfair comparisons between bot detectors.
+The issue arises from the inappropriate comparison of a general-purpose model with an extremely specialized—potentially overfitted—one.
+As an example, Hays et al.
+(2023) selected some datasets to train trivial classifiers with as few as four features.
+Then, they compared the trivial classifier to more complex, general, state-of-the-art models, on the same dataset where the trivial model had been trained, suggesting that very few features are sufficient to identify bots in a given dataset.
+While this approach can highlight biases in certain datasets, the comparison itself is unfair and should not be used to undermine the utility of general-purpose detectors or to criticize their performance.
+In addition to carefully choosing the models with which to compare performance, attention must also be devoted to the data used for the evaluation.
+As an example of problematic use of evaluation data, Gallwitz and Kreil (2022) selected prominent public accounts, such as members of the US Congress and Nobel Prize Laureates, to challenge the accuracy of a widely-used bot detection tool.
+These institutional accounts are usually managed by a team of social media managers via dedicated software, and not directly operated by the corresponding public figures.
+In other words, this dataset is a non-representative sample of the accounts that inhabit social platforms, and evaluating a general-purpose bot detector solely on these accounts likely leads to an under- or overestimation of the detector’s performance.
+Although relevant towards evaluating the performance of the detector in a very specific context, this evaluation does not warrant a generalized approval or criticism of the tool.
+More broadly, since bot detection methods—whether based on machine learning or on human annotation—have inherent error margins, accuracy results can be manipulated by carefully choosing test examples.
+In light of the pervasive challenges in evaluating results about social bots, a common recommendation is to always manually check a subset of the data after classification (Yang et al., 2022), as this often allows detecting outright misclassifications and possible underlying problems.
+Moreover, evaluating detectors should not only involve rigorous testing during training but also continuous validation when using pre-trained models, irrespective of their established reputation.
+Cherry-picking
+The issue of cherry-picking can manifest in unfair comparisons, but extends well beyond it.
 A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings.
-Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023) and to falsely claim that bot detectors have not been systematically evaluated (Gallwitz and Kreil, 2022).
+Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023).
 These claims can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020).
-On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—to give a false impression of prior research supporting their claims.
+On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—multiple times, to give a false impression of prior research supporting their claims.
 Such selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art in bot detection.
-By self-citing unvetted claims, authors create an illusion of authority.
+While citing unpublished work is not always bad per se, by self-citing under different forms multiple unvetted claims about the results of a single analysis, authors may create an illusion of authority.
 And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed to the same or comparable insights.
 This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of researchers who have previously advanced the field.
-Addressing the cherry-picking issue necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature.
+Addressing the cherry-picking issue in social bots research necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature, definitions, methods, or data.
 Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts.
 This discrepancy surfaces a crucial tension in maintaining the quality and rigor of peer review processes within rapidly evolving research landscapes (Van Noorden, 2023).
-The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of questionable quality.
+The field is now confronted not only with the difficulty of establishing a definitive benchmark, but also with the challenge of navigating through a heterogeneous literature landscape, where discerning genuine advancements becomes a complex task amid the noise produced by contributions of varying quality.
 Straw-man methodology
-The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation as if it faithfully represented the original research.
-For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on misrepresentation of the literature.
-As part of their criticism of supervised methods, they falsely claim that other researchers do not share or publish training and evaluation datasets.
-They also make demonstrably false claims about other researchers manipulating results through the use of unreported training data or threshold hacking.
-Gallwitz and Kreil criticize supervised bot detectors as invalid on the basis of these serious claims of scientific misconduct, yet do not provide any evidence in support of their allegation.
-Another methodological straw-man fallacy is the claim that Gallwitz and Kreil failed to reproduce an experiment by Keller and Klinger (2019).
-Gallwitz and Kreil applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in the bot detector.
-A different manifestation of the straw-man fallacy is the claim that bot detection is exclusively a supervised machine learning task.
-On the one hand, this assumption could be motivated by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of available supervised bot detectors (Cresci, 2020).
-On the other hand, such a view overlooks a similarly robust literature highlighting that supervised detectors suffer from lack of generalizability and transferability, as well as other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
-A variety of unsupervised approaches to bot detection have recently emerged as possible solutions to mitigate these problems (Cresci, 2020).
+The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation instead of the original research.
+For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on a misrepresentation of the literature.
+First, they claimed that «in our impression, most of the study authors we contacted were fully aware that the lion’s share or even all of the accounts they counted as ‘bots’ were, in fact, humans.» Such strong insinuation that the majority of researchers in the field knowingly misrepresented human accounts as bots is a serious allegation that requires robust evidence.
+However, this accusation is only based on the authors’ impression, lacking any scientific validity.
+Second, they claimed that a systematic evaluation of certain bot predictions in real-world scenarios have never been done prior to their work (Gallwitz and Kreil, 2022), pointing to a lack of manual validation and publicly available datasets for such evaluations.
+However, this claim is largely exaggerated.
+The majority existing bot detectors have been rigorously evaluated on training and test datasets, with many of these datasets being publicly available.
+While it is indeed important to manually validate the labels assigned by detectors and share this validation data, the generalized criticism that no systematic evaluation had been done before is unwarranted and potentially misleading, as such statement overlooks the extensive evaluations that have already been conducted in the field.
+Another problematic claim is that about Gallwitz and Kreil’s failure to reproduce an experiment by Keller and Klinger (2019).
+Here, they applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in Keller and Klinger’s detector, without mentioning that such differences could instead be due to the different choices they made with respect to Keller and Klinger.
+A further manifestation of the straw-man methodology is the claim that bot detection is exclusively a supervised machine learning task.
+This wrong assumption could perhaps be explained by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of existing supervised bot detectors (Cresci, 2020).
+Despite their prevalence, however, a robust body of literature highlighted that supervised detectors suffer from lack of generalizability and transferability, among other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
+A variety of unsupervised and semi-supervised bot detectors were proposed as possible solutions to these problems, considering groups instead of single accounts and studying features like the temporal patterns of their actions or the structures of their graph representations (Cresci, 2020).
 Evaluating the generalization capabilities of bot detectors is precisely the goal of Hays et al.
@@
 leverage their results to criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods.
-They do not make any distinction between the different types of bot detectors proposed to date—supervised ones based on the analysis of individual accounts vs.
+Similarly, Gallwitz and Kreil conclude that "the field of social bot research is fundamentally flawed", despite having investigated only a tiny fraction of the research based on one single bot detector.
+Neither of the two aforementioned works make any distinction between the different types of bot detectors proposed to date—including differences between supervised ones based on the analysis of individual accounts vs.
 unsupervised ones based on the analysis of groups of accounts (Cresci, 2020).
-This is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable.
+In (Hays et al., 2023), this is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable.
 However, the analysis only evaluates supervised detectors, which have known generalization issues (Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
 The unsurprising—and biased—result is that the considered detectors fail to generalize, which is used to criticize the whole field.
+Future work should avoid oversimplifying social bot detection as a solely supervised task.
+Instead, all works that propose new detectors, that use already existing ones, or that discuss the state-of-the-art, should acknowledge the multitude of approaches to the task, together with their strengths and weaknesses.
+Then, the type of bot detector to develop, use, or discuss in a given study should be chosen so as to be adequate and consistent with the objective of the study.
 Data biases
@@
 This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021).
-Some recent literature investigates the characteristics of accounts that are over ten years old.
+However, the majority of newly published bot detectors are at least partially based on dated—if not outright obsolete—training data.
 Obsolete datasets can be used to criticize other people’s detection methods or to promote one’s own method as better than the state of the art.
-A glaring example is the work of Hays et al.
-(2023), where out of eleven analyzed datasets, the most recent was published in 2020 and the least recent in 2011.
-Annotations can be another source of data bias.
-For example, so-called cyborg accounts (Ferrara et al., 2016) are automated in part via software, but closely managed and monitored by humans who can step in as needed to avoid detection.
-By labeling such accounts as human on the basis that they can reply to a query and assert that they are not bots, one can make it appear that a bot detection method has a higher false-positive rate than it actually does.
-To avoid this kind of bias, account labeling should be performed by multiple annotators following a shared and published rubric.
-Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) that accounts exhibit should be favored in place of binary labels.
-Some work in the literature fails to follow this methodology, only to accuse other researchers of bias.
-Gallwitz and Kreil (2022) demonstrate this behavior in several cases.
-Without disclosing any annotation rubric, they change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative.
+Annotations of accounts as humans or bots, or any other category, can introduce further bias when the annotation process is opaque or incoherent.
+As a practical example, cyborgs—accounts that are partly automated and partly managed by humans who can step in as needed to avoid detection—can be labeled as “humans” if one only considers their capacity to reply to a message.
+However, many bot detectors are based on the definition that any account with at least partial automation is a bot.
+When evaluating such bot detectors, labeling cyborgs as humans would unfairly inflate the false-positive rate due to the mismatch between the definition adopted by the bot detector versus that of the evaluator.
+To avoid this bias, account labeling should be performed by multiple independent annotators following a shared and openly accessible rubric, and based on a consistent definition of social bot.
+Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) should be favored in place of binary labels, as the former are better suited to capture the nuances in the use of automation exhibited by the multitude of online accounts.
+Unfortunately, some works in literature fail to follow these principles, resulting in inconsistent labeling and misleading results.
+Without disclosing any annotation rubric, Gallwitz and Kreil (2022) change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative.
 In one case, criticizing Yang et al.
-(2019), they label accounts that automatically post news headlines, such as those associated with major newswire agencies, as human.
-Their criticism of Keller and Klinger (2019) goes further: first, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger; second, they manually inspected only 0.04% of this different set; third, they ignored suspended accounts; and, fourth, they labeled as human accounts those that cross-posted tweets through software apps.
-Their conclusion was that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word.» Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al.
-(2020): suspended accounts are again assumed to be human to boost the alleged false-positive rate, then accounts that are clearly automated are still labeled as classification errors by changing the standard from “bot” to “malicious” or “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions is a dangerous trend toward scientific manipulation.
+(2019), they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as human.
+Their criticism of Keller and Klinger (2019) goes further.
+First, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger.
+Second, they manually inspected only 0.04% of this different set.
+Third, they ignored suspended accounts, and fourth, they labeled as human accounts those that cross-posted tweets through software apps.
+The aforementioned four arbitrary choices imply that they did not reproduce Keller and Klinger’s analysis.
+As such, their conclusion that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word» is simply false.
+Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al.
+(2020), as suspended accounts are again assumed to be human to boost the alleged false-positive rate.
+Then, accounts that are clearly automated are still presented as classification errors by labeling them as “(unmalicious) bot” and accounts “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions are among the malpractices that affect this field.
+Table 1: Summary of the main challenges and issues that may occur at each step of the social bots research process.
+For each issue, we propose a set of guidelines and practical recommendations.
+Since not all challenges apply to every study, it is up to the authors of future works to review this table to identify relevant issues specific to their study and follow the corresponding recommendations.
+The table could also serve as a valuable resource for reviewers and evaluators to assess the rigor of future works in terms of how new works implement the best practices and recommendations listed.
+step issues recommendations
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Practical recommendations
+The above sections highlighted a number of challenges and widespread methodological issues that currently affect the field of social bots research.
+Table LABEL:tab:checklist summarizes the main methodological issues and proposes a set of guidelines and practical recommendations to mitigate said issues so as to produce more reliable and robust research.
+The main issues and recommendations are also discussed in the following.
+One significant challenge is the use of contrasting definitions of social bots, which leads to incomparable results and confusion in research.
+To mitigate this, future studies should openly and clearly define the specific characteristics of the social bots they are targeting.
+Moreover, future research should adopt definitions consistently, without changing bot criteria and definitions case-by-case.
+Furthermore, favoring definitions based on objective characteristics—such as the use of automation—over subjective ones, will boost clarity and consistency.
+When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could have impacted the results.
+A related pressing issue is the use of inconsistent or opaque bot labeling procedures, as these hinder the clarity, comparability, and reproducibility of results.
+Researchers should document and publish their labeling procedures, including the criteria and tools used to carry out the annotation.
+Sharing labeled datasets with comprehensive metadata will further enhance transparency, reproducibility, and collaborative efforts.
+A further set of recommendations involve the evaluation procedures of social bot detectors.
+Newly trained bot detectors are often tested on well-known and old bots, which does not accurately reflect their performance in real-world scenarios.
+Researchers should prioritize using the most up-to-date datasets for their evaluations, as these are likely to better reflect the evolving nature of social bots.
+Performance assessments should include tests on bots with different and unknown characteristics, to better gauge the detectors’ real-world applicability and to ensure that detectors are robust and effective in dynamic online environments.
+In addition, attention to validation should not only be the focus of the training and testing phases of a new bot detector, but should also be devoted when using a pre-trained detector developed by others, even if it is well-known and established.
+A subset of automatically assigned labels and scores should always be manually validated, rather than accepted at face value.
+This ensures their accuracy and reliability in the given context, which might be different than that on which the detector was developed.
+As with the aforementioned manual labeling procedures, also this manual validation process should be clearly and openly documented, detailing the criteria and methods used for manual verification of the assigned labels.
+Another issue in the evaluation of bot detection performance is the possibility to cherry-pick competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector.
+To mitigate this issue, future research should benchmark bot detectors against a comprehensive set of existing detectors, including both strong and weak performers.
+Moreover, the criteria for the selection of the comparisons should be stated clearly and openly.
+Sharing evaluation scripts and datasets used in benchmarking will further promote transparency and reproducibility in performance comparisons.
+In addition, sensitivity analyses should also become a standard practice to address the possible sensitivity of detectors to thresholds and other model parameters, which are often overlooked when reporting bot detection results.
+These analyses should be reported alongside the main results to provide a clearer picture of the robustness and reliability of the detectors, aiding in the identification of optimal settings and limitations.
+Lastly, data availability and documentation are crucial for the progress of bot detection research.
+The lack of shared and well-documented datasets impedes reproducibility and comparison, and slows down the development of new detectors.
+Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labelled.
+This recommendation does not only involve the datasets used to train or evaluate a detector, such as those labeled manually, but also the datasets on which pre-trained detectors are applied.
+For the latter, it is crucial to share the data together with the labels or scores automatically assigned by the detectors.
+If data sharing is constrained by privacy or other issues, researchers should at least provide detailed descriptions of their datasets and the procedures for obtaining them.
+This practice will support reproducibility and facilitate more rigorous comparisons and advancements in bot detection research.
 (a) Simplistic and naïve model of social bots and human-operated accounts.
@@
 Figure 1: The ecology of online accounts: social bots and humans. Differences between naïve and realistic models, and their implications for data collection.
-Conceptual issues
+3 Conceptual issues
 Beyond methodological challenges, the landscape of social bots research is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues.
@@
 we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible.
-In addition to the many challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021; Torres-Lugo et al., 2022).
+In addition to the many general challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021).
 Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023).
@@
 This fundamental question cannot currently be answered.
+The previous discussion highlights the challenges in obtaining a reliable ground-truth for bot detection, which is largely due to the impossibility of acquiring accurate and dependable bot labels.
+Researchers are often compelled to use proxies, such as platform account removals or manual annotations based on profile characteristics, which however reduce the reliability and validity of the ground-truth.
+Furthermore, the lack of comprehensive knowledge about the types, sizes, and characteristics of bots operating on platforms creates an imbalance where data about certain bot types may be more accessible than others.
+Data availability can facilitate research on specific sub-types of bots, potentially leading to more favorable results, albeit confined to a limited subset of bots.
+However, this could also introduce bias in the development of bot detectors, making them particularly effective against the overrepresented bot types while being ineffective against the underrepresented ones.
 Common misconceptions in social bots research
-Part of the social bot literature claims that the bot detection task has been effectively solved.
-Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications.
-This optimistic view positions bot detection research as a success story for machine learning, highlighting the progress made in achieving practical and deployable solutions in the field.
+Some recent work has interpreted the results of social bot detection studies as a success story for machine learning, concluding that the positive results seem to suggest that the bot detection task is now effectively solved (Hays et al., 2023).
+Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions in the field.
 We disagree with this interpretation of the state of social bot research.
@@
 Below we report on some notable examples.
-The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019).
+The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019b).
 The limitations of benchmark datasets for bot detection were discussed by Feng et al.
@@
 [^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022).
-For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020; Pacheco et al., 2021).
+For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020).
 This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works.
@@
 Misconception 2 Bot detection performance can be improved easily.
-For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Cresci et al., 2021).
-Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term.
+For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020).
+Looking back at how this arms race has unfolded (Cresci et al., 2021), we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term.
 It is reasonable to assume that future advances will suffer a similar fate.
@@
 For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users.
-What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020; Pierri et al., 2023).
+What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020).
 Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022).
@@
 Misconception 6 All social bots research is flawed and bot detection results are useless.
-We see at least two strong arguments against this thesis.
+reference main result validation or corroboration source
+Yang et al. (2014) studied social bots and developed effective detection techniques platform collaborations in real-world scenarios platform deployments in production environments
+Subrahmanian et al.
+(2016) hosted and presented results of the DARPA Twitter bot challenge findings incorporated into defense and cybersecurity strategies U.S.
+Department of Defense reports and cybersecurity policies
+Ferrara et al.
+(2016) brought to light the issue of social bots and their capabilities in manipulation confirmed from various data leaks showing coordinated bot operations federal investigation reports, media disclosures
+Cresci et al.
+(2017) demonstrated that social bots evolve, becoming more sophisticated with time supported by subsequent studies and data showing similar patterns academic follow-up studies and industry research
+Varol et al. (2017) estimated social bots prevalence reports and admissions by social media platforms public statements and transparency reports from social media
+Lazer et al.
+(2018) highlighted the impact of bots on misinformation spread and public opinion alignment with findings from investigations into electoral interference government investigation documents and hearings
+Badawy et al. (2019) analyzed use of bots in Russian interference with the 2016 U.S. election corroborated by U.S. intelligence community findings U.S. Senate and House intelligence committee reports
+Cresci (2020) framed social bot detection as an adversarial machine learning task demonstrated by the improvements obtained via adversarial approaches academic follow-up studies and industry research
+Ferrara (2022) provided a taxonomy of Twitter bots, methods, and tools to detect them results supported by platforms and public organizations reports and investigations by government agencies
+Table 2: Examples of key findings on social bots that have withstood the test of time.
+Each result is accompanied by details on its validation or corroboration, along with the supporting sources.
+While not exhaustive, this list underscores the impact that rigorous social bots research can achieve.
+We see at least three strong arguments against this thesis.
 First, despite the limitations of bot detectors, there are several glaring examples of bot studies that were able to bring to light demonstrably harmful campaigns.
@@
 Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023).
-In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the findings.
+In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings.
 These cases are just some of the success stories of social bot research.
-If researchers had not developed bot detection techniques, we would not have been able to identify pockets of anomalous accounts engaged in sophisticated manipulations.
+Without the development of bot detection techniques, identifying clusters of anomalous accounts engaged in sophisticated manipulations would not have been possible.
 Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
-Second, the benefits of social bot research extend beyond the detection of malicious bots.
+Second, the methodological rigor and practical usefulness of some social bots studies are evidenced by corrective actions taken by government agencies based on academic research.
+For instance, in November 2014, the U.S.
+Securities and Exchange Commission (SEC) issued an alert to raise awareness about stock market manipulations exacerbated by social bots, [^7] Updated Investor Alert: Social Media and Investing – Avoiding Fraud (Nov.
+12, 2014) https://www.sec.gov/resources-for-investors/investor-alerts-bulletins/ia_socialmediafraud following the findings of earlier studies (Hwang et al., 2012).
+Subsequent research confirmed these suspicions in multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020a; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019).
+Both the 2018 and 2022 EU Code of Practice on Disinformation, as well as investigations by the U.S.
+Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee have also been informed by research on social bots.
+Furthermore, the concern over bot infiltration was highlighted when Elon Musk initially decided to abandon his purchase of Twitter, citing misleading information about the platform’s financial health and bot prevalence (Varol, 2023)—a claim also supported by Peiter Zatko, Twitter’s former head of security, who exposed serious deficiencies in the company’s security.
+[^8] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/ Table 2 provides some examples of contributions in the study and mitigation of social bots that have withstood the test of time.
+These examples underscore the enduring impact and significance of rigorous social bots research in shaping policy and industry practices.
+Third, the benefits of social bot research extend beyond the detection of malicious bots.
 For example, research and experimentation on social bots led to the development of neutral bots used to assess the level of political polarization and bias on a platform (Chen et al., 2021); “news bots” used for journalistic purposes to curate, aggregate, and distribute content gathered from multiple sources (Lokot and Diakopoulos, 2016); and even bots used for content moderation (Bilewicz et al., 2021; Askari et al., 2024).
 In addition, social bot research has contributed to the early development of other neighboring fields.
-The early research on social bots, which dates back to 2010 (Ratkiewicz et al., 2011; Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years.
+The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years.
 In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation.
@@
 In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
-Challenges in the post-API era
+4 Challenges in the post-API era
 In addition to the widespread biases and misconceptions detailed above, the field of social bot research confronts significant obstacles due to changes in social media platform policies.
-Previously, studies primarily concentrated on Twitter/X, likely because its free API facilitated data acquisition.
-However, this accessibility ceased in 2023 when Twitter/X terminated free data access for researchers.
+In the API era, most social bot studies have focused on Twitter/X due to its free API and the ease of large-scale data acquisition.
+This hyperfocus on Twitter/X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores platforms beyond Twitter/X.
+However, this data policy ceased in 2023 when Twitter/X terminated free data access for researchers.
 [^9] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram.
@@
 The advent of AI-powered social bots, which evade current bot detection models (Yang and Menczer, 2023; Yang et al., 2024), further emphasizes the importance of data availability in support of social bot research.
-Despite these setbacks, Europe’s Digital Service Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests.
+Despite these setbacks, Europe’s Digital Services Act (DSA) has offered some hope by mandating large social media platforms to grant researchers data access upon reasonable requests.
 [^11] https://digital-strategy.ec.europa.eu/en/policies/digital-services-act-package Platforms like TikTok, Meta, and Google have initiated new data access programs.
 However, these programs are significantly limited compared to previous access levels.
-Their opaque and stringent application review processes also leave their efficacy in question.
-Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics.
+Their opaque and stringent application review processes also leave their efficacy in question (Jaursch et al., 2024).
+Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduce new dynamics.
 Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors.
 The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community.
+Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth further contributes to hindering bot detection research on these new platforms.
 Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
-A call for moral responsibility
-Our analysis of the drawbacks and limitations of some recent studies brought to light the need for responsibility when discussing results about social bots.
+5 A call for moral responsibility
+Our analysis of the drawbacks and limitations of recent studies brought to light the need for responsibility when discussing results about social bots.
 The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion.
-To this end, it is paramount to avoid propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars—such as those between proponents and opponents of social bots research.
+To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars.
 The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society.
@@
 Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019).
-Indeed, we believe that all the literature that propagates the misconceptions demystified in this paper provides misleading contexts in some way.
+Perpetuating the methodological and conceptual issues addressed in this paper contributes to creating flawed or unreliable research.
 Cherry-picking data, references, claims, or results from a cited article are examples of misleading context in which misinformation thrives (West and Bergstrom, 2021).
@@
 This article aims to make a contribution in this direction by debunking common fallacious arguments adopted by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
-Acknowledgements
+6 Acknowledgments
 This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; by the European Union – Horizon 2020 Program under the scheme “INFRAIA-01-2018-2019 – Integrating Activities for Advanced Communities”, Grant Agreement n.

### v4

@@
 Overcoming such issues is instrumental towards ensuring reliable solutions and reaffirming the validity of the scientific method.
-In this contribution, we review some recent results in social bots research, highlighting and revising factual errors as well as methodological and conceptual biases.
+Here, we discuss a broad set of consequential methodological and conceptual issues that affect current social bots research, illustrating each with examples drawn from recent studies.
 More importantly, we demystify common misconceptions, addressing fundamental points on how social bots research is discussed.
 Our analysis surfaces the need to discuss research about online disinformation and manipulation in a rigorous, unbiased, and responsible way.
-This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research in the field.
+This article bolsters such effort by identifying and refuting common fallacious arguments used by both proponents and opponents of social bots research, as well as providing directions toward sound methodologies for future research.
 Corresponding author: stefano.cresci@iit.cnr.it
 1 Introduction
-Human decision-making processes depend on the availability of high-quality information.
-In addition, a healthy society requires a shared understanding of issues and values to foster trust and cooperation.
+Human decision-making processes depend on the availability of high-quality information and a healthy society requires a shared understanding of issues and values.
 Misinformation and disinformation erode trust and emphasize divisions.
-Moreover, the presence of divergent or incompatible beliefs can make it difficult to reach a consensus and take effective collective action.
-The repercussions could be so severe as to delay responses to a deadly pandemic (Pierri et al., 2022) or endanger critical democratic processes (Pennycook et al., 2021).
+Moreover, the presence of divergent or incompatible beliefs can hinder reaching consensus and taking effective collective action.
+The repercussions could be so severe as to delay responses to a deadly pandemic (Pierri et al., 2022) or endanger democratic processes (Pennycook et al., 2021).
 The science of misinformation seeks solutions to these problems (Lazer et al., 2018).
 However, this paramount endeavor can itself incur the same problems that it aims to overcome (West and Bergstrom, 2021; Altay et al., 2023).
-For example, scientific articles and publishers engage in a fierce competition for attention, similar to what is being experienced by news outlets.
-As a consequence, sensationalist claims and hyped results can sometimes be used as shortcuts to publication and other scientific accomplishments (West and Bergstrom, 2021).
-Then, confirmation bias is making its way into science in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021).
+For example, scientific articles and publishers engage in a fierce competition for attention, much like mainstream news outlets.
+As a consequence, sensationalist claims and hyped results are sometimes used as shortcuts to publication and scientific recognition (West and Bergstrom, 2021).
+Confirmation bias is also making its way into science, in the form of citation bias: the preference for citing articles that support one’s results over those that challenge them (West and Bergstrom, 2021).
 And again, over-generalizations of scientific results and poor understanding of methodological and conceptual limitations give rise to multiple misconceptions about misinformation (Altay et al., 2023).
@@
 The present study concerns one of the many forms of online manipulation: social bots.
-In this context, let us make a brief clarification of what, for the purposes of this article, is a bot, a social bot, and a malicious social bot.
-In literature, the term bot is associated with several others that are often, and sometimes erroneously, used as synonyms.
-As discussed by Gorwa and Guilbeault (2020b), scholars, journalists, and policymakers often speak indiscriminately of bots, chatbots, botnets, sybils, and cyborgs, to cite a few.
-If we go back in time, the bots better known as chatbots referred to early programs that can converse with humans.
-Later in time, the word bot also came to mean a computer machine compromised, co-opted, and remotely controlled by malware (Yang et al., 2014).
-When these types of accounts work together, they form a network called botnet, whose methods of operation and proposals for detection have been studied by information security researchers (Yang et al., 2014; Antonakakis et al., 2017).
-With the advent of social media and social networks, bots were also framed as compromised network accounts, called sybils in the information security field.
-Considering the bots that populate online social networks—conventionally called social bots (Ferrara et al., 2016)—there are works that categorize such accounts according their intent or their capacity to imitate humans (Stieglitz et al., 2018), as in the case of social bots promoting or suppressing certain political narratives (Pratelli et al., 2024) or facilitating the spread of disinformation (Shao et al., 2018).
-Others differentiated social bots according to their activity on the platform, identifying those that systematically retweet the posts of political candidates (Cresci et al., 2017), or those that mass-follow a brand or character to boost their popularity (Cresci et al., 2015).
-There even exist social bots with no malicious intent that rather serve a useful purpose for those who follow them (Cresci, 2020; Savage et al., 2016).
-For example, these bots can act as news aggregators (Lokot and Diakopoulos, 2016).
-Among the many diverse existing definitions, here we define bots as social media accounts that are totally or partially automated.
-By focusing on the use of automation, we adopt a definition that is both widely used and based on an objective criterion (Ferrara et al., 2016; Wischnewski et al., 2021; Ferrara, 2022).
-This is in contrast to those based on the intent (malicious, neutral, or benign), on the capacity to imitate humans, or on the type of activity carried out, which are inherently more subjective.
-Nonetheless, we clarify that independently of the chosen definition, the issues highlighted in the remainder of this article as well as the proposed solutions, are broadly valid and applicable, ensuring that adhering to a different definition does not diminish the overall validity of the analyses and discussions herein.
-Thanks to their automated nature, ease of use, and negligible marginal cost, social bots can be employed as expendable tools for spreading problematic content.
-Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb misinformation and disinformation (Shao et al., 2018).
+Among the many diverse existing definitions (Gorwa and Guilbeault, 2020), here we define social bots as social media accounts that are totally or partially automated.
+Hereafter, we use the terms “social bot” and “bot” interchangeably, always in line with the aforementioned definition.
+Due to their automation, simplicity, and low operating cost, social bots can be easily used as expendable tools for spreading problematic content.
+Understanding the role and activity of bots in large-scale manipulation campaigns is important, as it can inform strategies to curb mis- and disinformation (Shao et al., 2018).
 For this reason, social bots have attracted considerable scholarly and media attention (Allem and Ferrara, 2018; Assenmacher et al., 2020; González-Bailón and De Domenico, 2021; Chen et al., 2021).
-One of the latest such episodes was the public dispute on the real number of bots on Twitter that took place before Elon Musk’s acquisition of the platform (Varol, 2023).
+A recent example is the public dispute over Twitter’s bot count preceding Elon Musk’s acquisition (Varol, 2023).
 However, despite many years of research, the science of social bots is replete with the same problems that also plague the science of misinformation.
-These originate from uncorrected factual errors, from the presence of biases in how scientific results are cited and discussed, and from a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
-In our present analyses, we focus on some recently published articles about social bots.
-A thorough analysis of their methods, assumptions, results, and conclusions reveals several pitfalls ranging from factual inaccuracies and methodological flaws to more subtle incidents such as misrepresentation of the existing literature and conceptual errors.
-Our discussion links these issues to common misconceptions in social bots and misinformation research.
+These originate from uncorrected biases in how scientific results are cited and discussed, and a wide array of methodological and conceptual issues that set the stage for ambiguities, misunderstandings, unrealistic expectations, as well as contrasting and seemingly irreconcilable findings (Rauchfleisch and Kaiser, 2020; Hays et al., 2023).
+Here, we develop a critical and theoretically grounded perspective on key methodological and conceptual challenges in social bots research.
+To this end, we discuss a large set of consequential issues for the field, based on their recurrence, impact, and potential for misunderstanding.
+In discussing each issue, we refer to a small number of recent studies that serve as illustrative examples of the broader problems.
+Therefore, while this article does not aim to be exhaustive in terms of the analyzed literature, the range of issues addressed here is extensive and comprehensive—as reflected in Table LABEL:tab:checklist.
 Our critical analysis thus serves a twofold goal.
-On the one hand, we highlight and revise factual errors, methodological biases, and conceptual issues.
+On the one hand, we highlight and revise methodological biases and conceptual issues in social bots research.
 On the other hand, demystifying common misconceptions allows us to address fundamental issues on how science is produced and discussed (West and Bergstrom, 2021).
-Overcoming such issues is instrumental towards ensuring the credibility of science.
+Overcoming such issues is instrumental in ensuring the credibility of science.
 2 Methodological issues
-Social bot detection is a critical and increasingly challenging task in the realm of online safety and cybersecurity (Cresci, 2020).
-In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots—on various online platforms.
-The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts with the goal of detecting subtle differences between authentic users and automated entities.
+In this section, we focus on issues that arise from how social bots research is conducted and evaluated, covering aspects such as data selection, model training, and the dissemination of results.
+Social bot detection is a challenging task in the realm of online safety and cybersecurity (Cresci, 2020).
+In practice, it often involves the use of machine learning algorithms in a binary classification setting, with the goal of distinguishing human-operated accounts from automated ones—the social bots.
+The machine learning models are trained on features extracted from user profiles, behaviors, network interactions, and linguistic patterns acquired from user posts, with the goal of detecting subtle differences between authentic users and automated entities.
 Methodological issues in social bots research include a mix of typical machine learning pitfalls and unique issues inherent in the evolving nature of online social ecosystems, which we exemplify in the following sections.
 Model drift
-Model drift happens when a model relies on a set of features that sensibly diverge between train and test samples.
-Model drift generally happens when data slowly change distribution over time.
-However, it can also happen when the training and prediction datasets are significantly different.
-Generally, model drift can slowly reduce model accuracy over time, but it can also misleadingly show that the trained model performs well due to the exploitation of feature peculiarities not found in the prediction dataset.
+Model drift arises when the features a model relies on differ between training and prediction data—whether due to gradual shifts over time or inherent dataset differences.
+It can erode accuracy or falsely suggest strong performance by exploiting training-specific patterns absent at prediction time.
 In bot detection, the problem may surface when knowledge of some peculiar characteristic of the accounts is known beforehand and exploited at training time, allowing for near-perfect detection performance.
-For example, each botnet typically exhibits some peculiar characteristics that surface as a result of how the bots were created, their goals, or any other shared characteristic between them (Zhang et al., 2016; Mazza et al., 2022).
+For example, each botnet typically exhibits some peculiar characteristics resulting from the bot creation process, their goals, or any other shared characteristic (Zhang et al., 2016; Mazza et al., 2022).
 These peculiar characteristics might set the bots apart from the average behavior of human-operated accounts (Cresci et al., 2020).
-However, most of the time, such peculiar characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets.
-Here it is important to recall that the end goal of any bot detector—as with any machine learning classifier—is to accurately detect unknown bots from a set of never-seen-before data.
-Thus, the detection performance on a held-out portion of known data serves merely as a proxy for the real-world detection performance in-the-wild.
+However, typically such characteristics are known only for botnets that have already been identified and are instead unknown for still undetected botnets.
+Here we recall that a classifier’s detection performance on a held-out portion of known data serves merely as a proxy for its real-world detection performance in-the-wild.
+In fact, the end goal of any bot detector is to accurately detect unknown bots from a set of never-seen-before data.
 However, if a bot detector exploits known characteristics at training time for classifying instances of known bots, it might learn correlations that do not generalize to the unknown bots that the classifier will be tasked to detect later on.
-Another way in which a model drift occurs in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise).
+A specific instance of model drift in bot detection is when a detector exploits knowledge of how the accounts were collected—be them the malicious bots or the genuine ones—as a means to predict their class (i.e, automated or otherwise).
 The correlation between the information on how certain accounts were collected and their class is spurious with respect to the task of bot detection, which makes features based on such information dangerously misleading.
-Consequently, a detector exploiting these features would achieve excellent performance on known data but would likely exhibit poor capacity to generalize to new data with different characteristics.
-Such situation can lead to the proposal of ostensibly superior models, which however represent a regression rather than advancement in addressing the bot detection problem.
+Consequently, a detector exploiting these features would achieve excellent performance on known data but would exhibit poor capacity to generalize to new data with different characteristics.
+Leveraging features used to select and label the data can also be framed as a feature leakage issue, since those features would unfairly advantage the detector.
+Such a situation can lead to the proposal of ostensibly superior models, which, however, represent a regression rather than an advancement in addressing the bot detection problem.
 In addition, the same models could also be exploited to unduly undermine the performance of other models.
-This was recently done by Gallwitz and Kreil (2022) and Hays et al.
-(2023), who trained simple but unrealistic bot detectors using features tightly bound to the training dataset, such as whether an account was verified by Twitter—which was among the features used to select and label the data—to exaggerate the limitations of more general state-of-the-art models.
-The use of a feature used to select and label the data can also be framed as a feature leakage issue: it is crucial to refrain from using a feature that is utilized to select and label the data as a training input for the model since it will eventually lead to unreliable and inaccurate predictions.
+For example, some scholars trained simple but unrealistic bot detectors using features tightly bound to the training dataset, such as whether an account was verified by Twitter—which was among the features used to select and label the data—to exaggerate the limitations of more general state-of-the-art models (Gallwitz and Kreil, 2022; Hays et al., 2023).
 Unfair evaluation
-Another challenge in bot detection arises from the multitude of disparate definitions, detectors, and benchmark datasets (Cresci, 2020).
+Other challenges arise from the multitude of disparate definitions, detectors, and benchmark datasets (Cresci, 2020).
 For example, a critical issue in evaluating bot detectors is the possible inconsistency between the bot definition assumed by the detector and that used by the evaluator.
-This inconsistency can lead to biased and unfair evaluations as it is conceptually flawed to evaluate a detector using a different definition than that employed by the detector itself.
+This inconsistency can lead to biased and unfair evaluations, as it is conceptually flawed to evaluate a detector using a different definition than that employed by the detector itself.
 A tool trained to identify a specific type of bot should be assessed based on its ability to detect that specific type, not alternative ones for which it was not designed.
@@
 Therefore, while it is perfectly acceptable to discuss—and even criticize—the social bot definition used by a detector, evaluators must ensure that the definition used to evaluate a detector aligns with the detector’s, or be very clear about the differences and their implications towards the results of the evaluation.
-We point out, however, that there exists a trade-off between the need for consistency in using the same definition as the detector and the necessity of assessing the detector’s generalizability, which requires testing it against bots with different characteristics (Cresci, 2020).
+However, there exists an important trade-off between the need for consistency in using definitions and the necessity of assessing the detector’s generalizability, which requires testing it against bots with different characteristics (Cresci, 2020).
 Striking this balance depends on multiple factors, including the practical context in which the detector is used and whether the detector is specialized towards specific types of bots or general-purpose.
@@
 As an example, Hays et al.
-(2023) selected some datasets to train trivial classifiers with as few as four features.
-Then, they compared the trivial classifier to more complex, general, state-of-the-art models, on the same dataset where the trivial model had been trained, suggesting that very few features are sufficient to identify bots in a given dataset.
+(Hays et al., 2023) selected some datasets to train trivial classifiers with a very small number of features.
+Then, they compared the trivial classifiers to more complex, general, state-of-the-art models, on the same dataset where the trivial models had been trained, suggesting that very few features are sufficient to identify bots.
 While this approach can highlight biases in certain datasets, the comparison itself is unfair and should not be used to undermine the utility of general-purpose detectors or to criticize their performance.
-In addition to carefully choosing the models with which to compare performance, attention must also be devoted to the data used for the evaluation.
-As an example of problematic use of evaluation data, Gallwitz and Kreil (2022) selected prominent public accounts, such as members of the US Congress and Nobel Prize Laureates, to challenge the accuracy of a widely-used bot detection tool.
-These institutional accounts are usually managed by a team of social media managers via dedicated software, and not directly operated by the corresponding public figures.
-In other words, this dataset is a non-representative sample of the accounts that inhabit social platforms, and evaluating a general-purpose bot detector solely on these accounts likely leads to an under- or overestimation of the detector’s performance.
-Although relevant towards evaluating the performance of the detector in a very specific context, this evaluation does not warrant a generalized approval or criticism of the tool.
+Attention should also be devoted to the data used for the evaluation.
+As an example of problematic use of evaluation data, Gallwitz & Kreil (Gallwitz and Kreil, 2022) challenged the accuracy of a widely-used bot detector based on its results on a small dataset of public figures’ accounts.
+However, these accounts—typically managed by social media teams—do not represent the broader platform user base, possibly leading to under- or overestimation of the detector’s performance.
+While useful in a specific context, such evaluation does not justify generalized conclusions about the detector’s effectiveness.
 More broadly, since bot detection methods—whether based on machine learning or on human annotation—have inherent error margins, accuracy results can be manipulated by carefully choosing test examples.
@@
 Cherry-picking
-The issue of cherry-picking can manifest in unfair comparisons, but extends well beyond it.
+The issue of cherry-picking extends beyond unfair comparisons.
 A concerning trend involves selectively including, excluding, or misrepresenting prior literature to propose a narrative that aligns favorably with one’s own findings.
-Recently, this has been done to cast seemingly new criticism against some bot detectors (Hays et al., 2023).
-These claims can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020).
-On the other hand, Gallwitz and Kreil (2022) cite their own unpublished results—missing the vetting of peer review and shared via blog posts and presentations—multiple times, to give a false impression of prior research supporting their claims.
-Such selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art in bot detection.
+Recently, this has been done to cast seemingly new criticism against certain bot detectors (Hays et al., 2023).
+However, the impression of novelty in such criticism can only be made by omitting a significant amount of literature (Varol et al., 2017; Yang et al., 2019, 2020; Sayyadiharikandeh et al., 2020; Cresci, 2020).
+An opposite—but equally misleading—practice involves repeatedly citing one’s own unpublished results, which could give a false impression of prior research supporting one’s claims (Gallwitz and Kreil, 2022).
+Selective referencing can not only distort the perceived reliability and novelty of a work, but also skew the representation of the state-of-the-art.
 While citing unpublished work is not always bad per se, by self-citing under different forms multiple unvetted claims about the results of a single analysis, authors may create an illusion of authority.
-And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed to the same or comparable insights.
-This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of researchers who have previously advanced the field.
-Addressing the cherry-picking issue in social bots research necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature, definitions, methods, or data.
+And by omitting works that have already made certain contributions or conclusions, authors may create an illusion of originality, potentially overshadowing a substantial body of pre-existing research that has contributed comparable insights.
+This not only undermines the integrity of the academic discourse but also detracts from the collective acknowledgment and recognition owed to the broader community of scholars who have previously advanced the field.
+Addressing the cherry-picking issue necessitates a reliance on expert reviewers deeply versed in the nuances of the field, capable of discerning the strategic omission of relevant literature, definitions, methods, or data.
 Paradoxically, the escalating trend of publications (Haghani et al., 2022), particularly in hot topics like bot detection (Cresci, 2020), poses a formidable challenge, as the growing demand for reviewers outpaces the available pool of experts.
@@
 The straw-man fallacy consists of misrepresenting someone else’s research and then criticizing the misrepresentation instead of the original research.
-For example, Gallwitz and Kreil (2022) make a generic and sensationalist criticism against all social bots research based on a misrepresentation of the literature.
-First, they claimed that «in our impression, most of the study authors we contacted were fully aware that the lion’s share or even all of the accounts they counted as ‘bots’ were, in fact, humans.» Such strong insinuation that the majority of researchers in the field knowingly misrepresented human accounts as bots is a serious allegation that requires robust evidence.
-However, this accusation is only based on the authors’ impression, lacking any scientific validity.
-Second, they claimed that a systematic evaluation of certain bot predictions in real-world scenarios have never been done prior to their work (Gallwitz and Kreil, 2022), pointing to a lack of manual validation and publicly available datasets for such evaluations.
-However, this claim is largely exaggerated.
-The majority existing bot detectors have been rigorously evaluated on training and test datasets, with many of these datasets being publicly available.
-While it is indeed important to manually validate the labels assigned by detectors and share this validation data, the generalized criticism that no systematic evaluation had been done before is unwarranted and potentially misleading, as such statement overlooks the extensive evaluations that have already been conducted in the field.
-Another problematic claim is that about Gallwitz and Kreil’s failure to reproduce an experiment by Keller and Klinger (2019).
-Here, they applied a different bot detection model than the one used by Keller and Klinger, and then attributed the changes in bot score distributions to inconsistencies in Keller and Klinger’s detector, without mentioning that such differences could instead be due to the different choices they made with respect to Keller and Klinger.
-A further manifestation of the straw-man methodology is the claim that bot detection is exclusively a supervised machine learning task.
+A common manifestation is the claim that bot detection is exclusively a supervised machine learning task.
 This wrong assumption could perhaps be explained by the fact that supervised machine learning is the traditional way in which the task was tackled, and by the multitude of existing supervised bot detectors (Cresci, 2020).
-Despite their prevalence, however, a robust body of literature highlighted that supervised detectors suffer from lack of generalizability and transferability, among other issues (Cresci, 2020; Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
-A variety of unsupervised and semi-supervised bot detectors were proposed as possible solutions to these problems, considering groups instead of single accounts and studying features like the temporal patterns of their actions or the structures of their graph representations (Cresci, 2020).
-Evaluating the generalization capabilities of bot detectors is precisely the goal of Hays et al.
-(2023).
-However, they chose to do so by adopting a supervised learning methodology that has been proven to suffer from critical issues when it comes to generalizability and transferability.
-Another issue with their methodology is that it does not fully support their conclusions.
-Specifically, Hays et al.
-leverage their results to criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods.
-Similarly, Gallwitz and Kreil conclude that "the field of social bot research is fundamentally flawed", despite having investigated only a tiny fraction of the research based on one single bot detector.
-Neither of the two aforementioned works make any distinction between the different types of bot detectors proposed to date—including differences between supervised ones based on the analysis of individual accounts vs.
-unsupervised ones based on the analysis of groups of accounts (Cresci, 2020).
-In (Hays et al., 2023), this is a textbook example of circular reasoning: the goal of the analysis is to evaluate whether current bot detectors are generalizable.
-However, the analysis only evaluates supervised detectors, which have known generalization issues (Echeverría et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
-The unsurprising—and biased—result is that the considered detectors fail to generalize, which is used to criticize the whole field.
-Future work should avoid oversimplifying social bot detection as a solely supervised task.
-Instead, all works that propose new detectors, that use already existing ones, or that discuss the state-of-the-art, should acknowledge the multitude of approaches to the task, together with their strengths and weaknesses.
-Then, the type of bot detector to develop, use, or discuss in a given study should be chosen so as to be adequate and consistent with the objective of the study.
+Despite their prevalence, however, a robust body of literature highlighted that supervised detectors suffer from lack of generalizability and transferability (Cresci, 2020; Echeverr´ıa et al., 2018; Yang et al., 2020; Sayyadiharikandeh et al., 2020; Rauchfleisch and Kaiser, 2020; Dimitriadis et al., 2021).
+A variety of unsupervised and semi-supervised detectors were proposed as possible solutions to these problems, considering groups instead of single accounts and studying features like the temporal patterns of their actions or the structures of their graph representations (Cresci, 2020).
+Nonetheless, some studies specifically focus on evaluating the generalization capabilities of bot detectors, but only consider supervised detectors with known generalization deficiencies.
+The unsurprising result is that the considered detectors fail to generalize, which is used to criticize the whole field—a textbook example of circular reasoning.
+For instance, Hays et al.
+(Hays et al., 2023) criticize the generalization capabilities of all bot detectors, despite having experimented only with a narrow set of supervised methods.
+Similarly, Gallwitz & Kreil (Gallwitz and Kreil, 2022) conclude that «the field of social bot research is fundamentally flawed», despite having investigated only a tiny fraction of the research based on one single bot detector.
+The issue lies in the broad conclusions of such studies, which are unsupported by their methodology and results.
+Neither of the two aforementioned works made distinctions between the different types of bot detectors proposed to date.
+A further example is the claim that a systematic evaluation of bot predictions in real-world scenarios have never been done (Gallwitz and Kreil, 2022).
+The accusation points to a lack of manual validation and publicly available datasets.
+However, the majority of existing bot detectors have indeed been evaluated on training and test datasets, with many of these datasets being publicly available.
+While we acknowledge that such evaluations may have inherent limitations, and that it is important to manually validate bot labels and to share this validation data, the generic criticism that no systematic evaluation has been done before is unwarranted and potentially misleading.
+Therefore, future work should avoid oversimplifying social bot detection as a solely supervised task.
+Instead, works that propose new detectors, that use already existing ones, or that discuss the state-of-the-art, should acknowledge the multitude of approaches to the task, with their strengths and weaknesses.
+Then, the type of evaluation, as well as the type of bot detector to develop, use, or discuss in a given study should be chosen so as to be adequate and consistent with the objective of the study.
 Data biases
-In machine learning and data science tasks, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009).
+In machine learning and data science, high-quality data is the linchpin for robust model development and insightful analyses (Halevy et al., 2009).
 As mentioned above, numerous datasets of bot and human accounts have been published over the years.
@@
 When considering the representativeness of published benchmark datasets with respect to the bots that currently inhabit online platforms, the readers should be aware that the accounts included in a dataset are likely a few years older than the publication date of the dataset itself.
-Moreover, some datasets are built by combining or extending previously published resources.
-For example, the pan-2019 dataset (Rangel and Rosso, 2019) is largely based on other datasets published between 2011 and 2017.
 Hence, the publication date of a dataset should be considered as a generous upper bound of the recency of the accounts therein.
-Even under this relaxed assumption, in consideration of the rapid pace of change of the landscape of online harms and of the availability of more comprehensive and more recent datasets (Feng et al., 2021), we may run the risk of capturing datasets that are no longer representative of the actual state of the platforms.
+Even under this relaxed assumption, in consideration of the rapidly evolving landscape of online harms and the availability of more comprehensive and recent datasets (Feng et al., 2021), we risk relying on datasets that are no longer representative of the actual state of the platforms.
 This is particularly troublesome in light of the known evolutionary behavior of social bots, typical of adversarial settings, which requires constant updates of data and methods (Cresci et al., 2021).
-However, the majority of newly published bot detectors are at least partially based on dated—if not outright obsolete—training data.
-Obsolete datasets can be used to criticize other people’s detection methods or to promote one’s own method as better than the state of the art.
+However, many newly published bot detectors are at least partially based on obsolete training data.
 Annotations of accounts as humans or bots, or any other category, can introduce further bias when the annotation process is opaque or incoherent.
 As a practical example, cyborgs—accounts that are partly automated and partly managed by humans who can step in as needed to avoid detection—can be labeled as “humans” if one only considers their capacity to reply to a message.
-However, many bot detectors are based on the definition that any account with at least partial automation is a bot.
+However, many bot detectors are based on the definition that any account with partial automation is a bot.
 When evaluating such bot detectors, labeling cyborgs as humans would unfairly inflate the false-positive rate due to the mismatch between the definition adopted by the bot detector versus that of the evaluator.
-To avoid this bias, account labeling should be performed by multiple independent annotators following a shared and openly accessible rubric, and based on a consistent definition of social bot.
-Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) should be favored in place of binary labels, as the former are better suited to capture the nuances in the use of automation exhibited by the multitude of online accounts.
-Unfortunately, some works in literature fail to follow these principles, resulting in inconsistent labeling and misleading results.
-Without disclosing any annotation rubric, Gallwitz and Kreil (2022) change criteria for labeling bots multiple times, thus inflating the false-positive or false-negative rates of bot detection methods to fit their narrative.
-In one case, criticizing Yang et al.
-(2019), they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as human.
-Their criticism of Keller and Klinger (2019) goes further.
-First, Gallwitz and Kreil considered a different set of accounts than those analyzed by Keller and Klinger.
-Second, they manually inspected only 0.04% of this different set.
-Third, they ignored suspended accounts, and fourth, they labeled as human accounts those that cross-posted tweets through software apps.
-The aforementioned four arbitrary choices imply that they did not reproduce Keller and Klinger’s analysis.
-As such, their conclusion that «not a single one of the accounts in [Keller and Klinger’s] list was a “bot” in any meaningful sense of the word» is simply false.
-Similarly biased annotations are used by Gallwitz and Kreil (2022) to criticize Dunn et al.
-(2020), as suspended accounts are again assumed to be human to boost the alleged false-positive rate.
-Then, accounts that are clearly automated are still presented as classification errors by labeling them as “(unmalicious) bot” and accounts “spreading vaccine-critical information.” The use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions are among the malpractices that affect this field.
+In other cases, the authors change criteria for labeling bots multiple times within the context of the same study, without disclosing any annotation rubric (Gallwitz and Kreil, 2022).
+First, they label accounts that automatically post news headlines—such as those associated with major newswire agencies—as humans.
+Then, they ignore suspended accounts or consider them to be human, depending on the analysis.
+Finally, they label accounts that cross-post tweets through software apps as humans.
+As exemplified above, the use of obsolete or biased datasets with unclear, inconsistent, or shifting definitions and opaque labeling schemes are among the malpractices that affect this field.
+To avoid bias, account labeling should be performed by multiple independent annotators following a shared and openly accessible rubric, and based on a consistent definition of social bot.
+Moreover, continuous scores representing the different degrees of “botness” (i.e., automation) should be favored in place of binary labels, as the former are better suited to capture nuances in the use of automation.
 Table 1: Summary of the main challenges and issues that may occur at each step of the social bots research process.
@@
 Practical recommendations
-The above sections highlighted a number of challenges and widespread methodological issues that currently affect the field of social bots research.
-Table LABEL:tab:checklist summarizes the main methodological issues and proposes a set of guidelines and practical recommendations to mitigate said issues so as to produce more reliable and robust research.
-The main issues and recommendations are also discussed in the following.
+Table LABEL:tab:checklist summarizes the main methodological issues that may occur at each step of the social bots research process, and proposes a set of best practices, guidelines, and practical recommendations to mitigate them.
 One significant challenge is the use of contrasting definitions of social bots, which leads to incomparable results and confusion in research.
@@
 Furthermore, favoring definitions based on objective characteristics—such as the use of automation—over subjective ones, will boost clarity and consistency.
-When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could have impacted the results.
+When discussing results, researchers should also compare their definitions with existing ones to highlight similarities and differences, also exploring how different definitions could impact the results.
 A related pressing issue is the use of inconsistent or opaque bot labeling procedures, as these hinder the clarity, comparability, and reproducibility of results.
@@
 Sharing labeled datasets with comprehensive metadata will further enhance transparency, reproducibility, and collaborative efforts.
-A further set of recommendations involve the evaluation procedures of social bot detectors.
+A further set of recommendations involves the evaluation procedures of social bot detectors.
 Newly trained bot detectors are often tested on well-known and old bots, which does not accurately reflect their performance in real-world scenarios.
@@
 As with the aforementioned manual labeling procedures, also this manual validation process should be clearly and openly documented, detailing the criteria and methods used for manual verification of the assigned labels.
-Another issue in the evaluation of bot detection performance is the possibility to cherry-pick competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector.
+Another issue in the evaluation of bot detection performance is the possibility of cherry-picking competitors in such a way to purportedly demonstrate the superior performance of a newly proposed detector.
 To mitigate this issue, future research should benchmark bot detectors against a comprehensive set of existing detectors, including both strong and weak performers.
@@
 The lack of shared and well-documented datasets impedes reproducibility and comparison, and slows down the development of new detectors.
-Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labelled.
+Researchers should make their datasets publicly available whenever possible, ensuring they are well-documented with comprehensive metadata, including how the data was collected and labeled.
 This recommendation does not only involve the datasets used to train or evaluate a detector, such as those labeled manually, but also the datasets on which pre-trained detectors are applied.
@@
 This practice will support reproducibility and facilitate more rigorous comparisons and advancements in bot detection research.
-(a) Simplistic and naïve model of social bots and human-operated accounts.
-Here, all bots are alike, they are evenly distributed, and we have complete knowledge on their numbers and characteristics.
+(a) Simplistic and naïve model of social bots (black-colored) and human-operated accounts (grey-colored).
+Here, all bots are alike, they are evenly distributed, and we have complete knowledge of their numbers and characteristics.
 According to this model, collecting unbiased and comprehensive bot datasets is possible.
@@
 3 Conceptual issues
-Beyond methodological challenges, the landscape of social bots research is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues.
+Beyond methodological challenges, the field is also intricately intertwined with a variety of potentially more subtle—yet profoundly influential—conceptual issues.
+These concern how social bots and related phenomena are defined, framed, and understood within the research discourse.
 These nuanced challenges, if left unaddressed, have the potential to exert an even more insidious influence on the development of the field and its reception.
 Failure to account for context
-In discussing the limitations that currently hinder progress in bot detection, some recent research has pointed the finger at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023).
-This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in datasets acquired with rigorous practices, contrary to the literature so far.
-More specifically, this criticism of the existing data collection practices subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient.
+In discussing the limitations that currently hinder progress in bot detection, some recent research pointed at flawed data collection practices that fall short of capturing the complexity of the bot space (Hays et al., 2023).
+This interpretation of the limitations of the existing datasets implicitly assumes the possibility to encode the full complexity of the bot detection problem space in a dataset.
+More specifically, this criticism subsumes that building unbiased and comprehensive bot datasets is possible, and perhaps even convenient.
 Unbiased datasets are needed so as to avoid that peculiar bot and human characteristics leak into the data allowing simple models to achieve good performance on the collected data, while showing poor performance in real scenarios.
@@
 Regarding bias, many bot datasets were found to contain biases, in that the accounts therein featured some peculiar characteristics (Hays et al., 2023).
-We note that this is largely unsurprising since such peculiar characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a single entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016).
-Given that all bots belonging to the same botnet are created and/or operated by the same entity, and pursue the same common goal, they tend to share some similarities.
+However, this is largely unsurprising since such characteristics are the inevitable consequence of the bots being organized in botnets: groups of accounts created and controlled by a central entity who collectively operates the bots to reach some predefined goal (Zhang et al., 2016).
+Given that all bots belonging to the same botnet are created or operated by the same entity, and pursue the same common goal, they tend to share similarities.
 When compared to other accounts, the botnet thus appears to have some peculiar characteristics, as highlighted in Figure 1(b).
-Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather should be traced back to the very nature of the social bot phenomenon.
+Thus, in most cases, the presence of biases in the existing bot datasets should not be blamed on the creators of the datasets, but rather traced back to the very nature of the social bot phenomenon.
 The assumption about the possibility to build comprehensive bot datasets is equally flawed.
@@
 This would allow obtaining an accurate representation of the bot landscape that captures the full complexity of the problem space.
-While in some domains of application for machine learning this is feasible, [^2] For example, in certain speech recognition tasks all instances (e.g., languages) to be classified are known in advance.
-Given a sufficiently large and diverse set of user inputs, unbiased and comprehensive datasets are possible (Ardila et al., 2020).
-we argue that performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible.
-In addition to the many general challenges brought by social data (Olteanu et al., 2019; Pasquetto et al., 2020), the task of detecting social bots also partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021).
-Indeed, there has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023).
+However, performing a uniform and random sample of the whole population of bots is extremely problematic—if not outright impossible.
+The task of detecting social bots partly belongs to the fields of information security and open-source intelligence, which are intrinsically characterized by the presence of adversaries who are strongly motivated to remain hidden (Cresci et al., 2021).
+There has always been limited knowledge of the real extent of the bot problem, which makes it hard to track bots on a platform, quantify their numbers, and assess their impact (Varol, 2023; Mendoza et al., 2020; Tan et al., 2023).
 We only have a partial understanding and visibility of the botnets that operate on our online platforms.
@@
 How can we create unbiased and comprehensive datasets when we largely ignore the types and numbers of bots populating online platforms?
-This fundamental question cannot currently be answered.
-The previous discussion highlights the challenges in obtaining a reliable ground-truth for bot detection, which is largely due to the impossibility of acquiring accurate and dependable bot labels.
-Researchers are often compelled to use proxies, such as platform account removals or manual annotations based on profile characteristics, which however reduce the reliability and validity of the ground-truth.
-Furthermore, the lack of comprehensive knowledge about the types, sizes, and characteristics of bots operating on platforms creates an imbalance where data about certain bot types may be more accessible than others.
-Data availability can facilitate research on specific sub-types of bots, potentially leading to more favorable results, albeit confined to a limited subset of bots.
-However, this could also introduce bias in the development of bot detectors, making them particularly effective against the overrepresented bot types while being ineffective against the underrepresented ones.
+This fundamental question is currently unanswered.
 Common misconceptions in social bots research
-Some recent work has interpreted the results of social bot detection studies as a success story for machine learning, concluding that the positive results seem to suggest that the bot detection task is now effectively solved (Hays et al., 2023).
-Proponents of this perspective highlight the perceived accuracy and reliability of current detection methods, and advocate their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions in the field.
-We disagree with this interpretation of the state of social bot research.
-On the contrary, we believe that unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection.
-We are well aware of this body of work, having actively contributed to its development for many years.
-Below we report on some notable examples.
+Some recent work posed the idea that the seemingly excellent results in social bot detection could be interpreted as a success story for machine learning and as evidence that the task is now effectively solved (Hays et al., 2023).
+According to this perspective, the perceived accuracy and reliability of current detection methods support their suitability for integration into various downstream applications, highlighting the progress made in achieving practical and deployable solutions.
+While Hays et al.
+(Hays et al., 2023) pose the question “Is bot detection a solved problem?” as a rhetorical device, both their framing and conclusions imply that the field has collectively overlooked critical challenges.
+This is a problematic foundation.
+A vast literature has long acknowledged the limitations in social bots research.
+By sidestepping this body of work, the authors construct a straw man: a field that blindly celebrates near-perfect performance without reflection.
+Other than boosting the apparent novelty of their critique, this framing misrepresents the field.
+Unbiased and responsible research in this area should acknowledge the body of literature warning researchers and practitioners about the limitations of bot detection.
+Below, we report on some notable examples.
 The first accounts of the difficulties supervised detectors face in detecting sophisticated bots date back to 2013 (Yang et al., 2013) and intensified in subsequent years (Cresci et al., 2017; Grimme et al., 2018; Cresci et al., 2019b).
 The limitations of benchmark datasets for bot detection were discussed by Feng et al.
-(2021) and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al.
-(2022).
-Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverría et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021).
+(Feng et al., 2021), and the data-related limitations to the generalizability and replicability in social bots research were touched upon by Assenmacher et al.
+(Assenmacher et al., 2022).
+Several works measured and discussed the generalization deficiencies of bot detectors, with particular emphasis on supervised ones (Echeverr´ıa et al., 2018; Yang et al., 2019; Cresci, 2020; Sayyadiharikandeh et al., 2020; Yang et al., 2020; Feng et al., 2021).
 Others warned about the lack of consistency in bot classifications obtained at different points in time (Rauchfleisch and Kaiser, 2020) and with different detectors (Svenaeus, 2020; Martini et al., 2021).
-Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), as well as by independent researchers and the media.
-[^4] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^5] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^6] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the continuing challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022).
-For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated malicious behavior (Cresci, 2020; Roth and Pickles, 2020).
+Concerns about the accuracy and usefulness of bot detection have also been raised by security and integrity experts at major platforms (Roth and Pickles, 2020), and by independent researchers and the media.
+[^3] https://www.nytimes.com/2020/06/16/science/social-media-bots-kazemi.html [^4] https://mediawell.ssrc.org/expert-reflections/on-digital-disinformation-and-democratic-myths/ [^5] https://www.lawfareblog.com/random-toxicity-whats-going-benjaminwittess-mentions The growing tension between the lasting challenges of bot detection and the increasing technological capabilities of online manipulators, including unprecedented affordances such as those offered by generative AI (Yang and Menczer, 2023; Ferrara, 2023), even led some researchers to question the long-term viability of the bot detection task (Boneh et al., 2019; Grimme et al., 2022).
+For the same reasons, some have suggested redirecting part of the ongoing scientific effort away from bot detection and to alternative and more promising tasks, such as detecting information operations and coordinated harmful behavior (Cresci, 2020; Roth and Pickles, 2020; Mannocci et al., 2024).
 This sampling of the existing literature on the challenges and limitations of bot detection paints a very different picture than some recent, optimistic works.
@@
 Misconception 1 Social bot detection is a solved task.
-The above evidence shows that despite significant efforts devoted for a prolonged time (Cresci, 2020), bot detection is nowhere near to being a solved problem—quite the contrary.
+The above evidence shows that despite significant efforts devoted for a prolonged time, bot detection is nowhere near to being a solved problem—quite the contrary.
 In framing the bot detection task as a seemingly straightforward problem, some authors also hint at the possibility of easily adapting existing detectors to overcome their current limitations and keep pace with the evolution of more human-like bots (Hays et al., 2023).
@@
 Misconception 2 Bot detection performance can be improved easily.
-For many years, we have witnessed a game of whack-a-mole between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020).
-Looking back at how this arms race has unfolded (Cresci et al., 2021), we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term.
+For many years, we witnessed to a whack-a-mole game between bot developers, with their increasingly sophisticated accounts (Yang and Menczer, 2023; Yang et al., 2024), and bot hunters, equipped with a variety of different detectors (Yang et al., 2019; Sayyadiharikandeh et al., 2020; Yang et al., 2020).
+Looking back at how this arms race has unfolded, we can conclude that none of the technological advances we have experimented with have significantly mitigated the challenges posed by malicious social bots in the long term.
 It is reasonable to assume that future advances will suffer a similar fate.
-Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology.
+Social bots are adversarial, fast-moving targets, characterized by a fast adoption of cutting-edge technology (Cresci et al., 2021).
 Bot detection is therefore an intrinsically challenging task, made even more daunting by the lack of accurate information on the targets of the analysis, the limited collaboration from online platforms, and the rapidly evolving nature of online harms.
-Below we discuss some other misconceptions that undermine our understanding of social bots and of the related literature:
+Below, we discuss some other misconceptions that undermine the understanding of social bots and of the related literature:
 Misconception 3 All social bots are similar.
 Misconception 4 Each bot detector can detect all types of bots.
-A common assumption in the scientific literature, as well as among the general public, is that if a given bot detector has achieved good detection performance against some bots, then that detector should be able to replicate that performance on any bot detection task.
-This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022) and only considering the known botnets.
+Both scholars and the general public often assume that a bot detector performing well on some bots will perform equally well on any bot detection task.
+This assumption is problematic given that there are a plethora of diverse bots, each with its own characteristics (Mazza et al., 2022).
 Consider, for example, the differences between the bots used to boost the popularity of certain public figures—so-called fake followers (Cresci et al., 2015)—and those that manipulate trending topics—astroturfers and spammers (Abokhodair et al., 2015).
 Or the differences between bots involved in political manipulation (Shao et al., 2018; Caldarelli et al., 2020) versus those aimed at fooling automated trading systems (Tardelli et al., 2022; Bello et al., 2023).
-In other words, not all bots are the same.
 Some scholars addressed this heterogeneity by designing detectors that aim for generality and broad applicability, such as Botometer, especially in its latest releases (Sayyadiharikandeh et al., 2020; Yang et al., 2020).
 Others have developed detectors specifically designed to detect certain types of bots.
-The latter trade generalizability and portability for detection accuracy.
-Their performance depends heavily on the characteristics of the bots they are designed to detect.
+The latter trade generalizability and portability for detection accuracy, and their performance depends heavily on the characteristics of the bots they are designed to detect.
 For example, a detector designed to detect time-synchronized retweeting actions (Mazza et al., 2019) would likely be useless at detecting mass-following bots (Cresci et al., 2015).
 However, this should not be regarded as a limitation of the bot detector—let alone one deliberately unstated by its developers in order to pass it off as a good product—but rather as an inappropriate use of the detector itself.
-Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized detectors.
-In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot or not.
+Limitations in generalizability and portability also affect general-purpose bot detectors, although to a lesser extent than specialized ones.
+In fact, even general-purpose detectors rely on a limited number of features to estimate whether an account is a bot.
 Therefore, in general, any detector, specialized or otherwise, has variable performance and its detection capabilities depend on the characteristics of the accounts it is applied to.
 In conclusion, no single bot detector is capable of detecting all types of bots.
-Many articles proposing a new bot detector, or investigating bot interference, motivate the work on the basis of the role of social bots in spreading disinformation (Shao et al., 2018).
-Similar arguments are made for the role of bots in exacerbating social conflicts (Stella et al., 2018).
-In other words, part of the social bot literature is based on the assumption that bots are responsible for many of the problems that affect our online platforms and the information ecosystem as a whole.
-This brings to the fore another misconception:
+Social bots research is often framed within the broader landscape of mis- and disinformation studies, a field that extends well beyond online manipulation, including critical perspectives on post-truth dynamics, cognitive biases, epistemic trust, ideological polarization, and the structural features of media ecosystems (Spohr, 2017; Waisbord, 2018).
+The specific relevance of social bots in this broader discourse has largely emerged from concerns over their potential to enable scalable and low-cost manipulation (Shao et al., 2018; Stella et al., 2018).
+These concerns, while widely shared and influential, are themselves being increasingly reexamined, raising questions about the actual impact of bots, the assumptions behind their perceived threat, and the relative emphasis they have received in disinformation research.
+This surfaces another misconception:
 Misconception 5 Social bots are mainly responsible for the spread of disinformation.
-An unbiased analysis of the existing literature paints a much more dubious picture of the role of social bots.
+An unbiased analysis of the existing literature paints a dubious picture of the role of social bots.
 For example, while some studies have concluded that bots play a prominent role in the spread of problematic content (Shao et al., 2018; Stella et al., 2018), others have reached the opposite conclusion (Vosoughi et al., 2018; González-Bailón and De Domenico, 2021; Seckin et al., 2024).
@@
 For these reasons, we currently lack scientific consensus and conclusive evidence about the role of social bots and their effectiveness in influencing online users.
-What we do know, however, is that bots are only one of many agents involved in the spread of misinformation and disinformation (Starbird, 2019; Roth and Pickles, 2020).
-Examples of some of the other agents include state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022).
+What we do know, however, is that bots are only one of many agents involved in the spread of mis- and disinformation (Starbird, 2019; Roth and Pickles, 2020).
+Examples of other agents are state-sponsored trolls, users who collude and coordinate for malicious purposes, superspreaders, and even willing but unwitting individuals (Starbird, 2019; DeVerna et al., 2022).
 Each of these agents represents a potential threat to safe and trusted online platforms, and a thriving area of research and experimentation.
-It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overusing some while overlooking others, based on unsupported decisions.
+It is therefore critical to balance efforts in all of these directions, avoiding the pitfall of overstudying some while overlooking others, based on unsupported decisions.
 The previous misconceptions might induce readers to think that social bot research has led to flawed—if not outright useless—results.
-Such a claim has been made by Hays et al.
-(2023).
-Similarly, Gallwitz and Kreil (2022) not only question the validity of existing bot detection methods—based on the fallacies discussed earlier—but go even further in arguing that social bots are not an issue.
+Such claims have been recently made by Hays et al.
+(Hays et al., 2023) and Gallwitz & Kreil (Gallwitz and Kreil, 2022)—the latter even in sensationalist terms.
 These claims lead to our last, but by no means least important, misconception:
 Misconception 6 All social bots research is flawed and bot detection results are useless.
-reference main result validation or corroboration source
-Yang et al. (2014) studied social bots and developed effective detection techniques platform collaborations in real-world scenarios platform deployments in production environments
-Subrahmanian et al.
-(2016) hosted and presented results of the DARPA Twitter bot challenge findings incorporated into defense and cybersecurity strategies U.S.
-Department of Defense reports and cybersecurity policies
-Ferrara et al.
-(2016) brought to light the issue of social bots and their capabilities in manipulation confirmed from various data leaks showing coordinated bot operations federal investigation reports, media disclosures
-Cresci et al.
-(2017) demonstrated that social bots evolve, becoming more sophisticated with time supported by subsequent studies and data showing similar patterns academic follow-up studies and industry research
-Varol et al. (2017) estimated social bots prevalence reports and admissions by social media platforms public statements and transparency reports from social media
-Lazer et al.
-(2018) highlighted the impact of bots on misinformation spread and public opinion alignment with findings from investigations into electoral interference government investigation documents and hearings
-Badawy et al. (2019) analyzed use of bots in Russian interference with the 2016 U.S. election corroborated by U.S. intelligence community findings U.S. Senate and House intelligence committee reports
-Cresci (2020) framed social bot detection as an adversarial machine learning task demonstrated by the improvements obtained via adversarial approaches academic follow-up studies and industry research
-Ferrara (2022) provided a taxonomy of Twitter bots, methods, and tools to detect them results supported by platforms and public organizations reports and investigations by government agencies
-Table 2: Examples of key findings on social bots that have withstood the test of time.
-Each result is accompanied by details on its validation or corroboration, along with the supporting sources.
-While not exhaustive, this list underscores the impact that rigorous social bots research can achieve.
 We see at least three strong arguments against this thesis.
@@
 For example, detectors developed as part of some scientific efforts were later deployed on online platforms and used to remove large numbers of malicious accounts (Yang et al., 2014).
-Similarly, the results of some bot activity studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023).
-In several other cases, scientific findings about bot activity were later found to be largely consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings.
-These cases are just some of the success stories of social bot research.
-Without the development of bot detection techniques, identifying clusters of anomalous accounts engaged in sophisticated manipulations would not have been possible.
-Therefore, even if no universal bot detector exists yet, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
+Similarly, the results of some studies led platforms to remove accounts identified as malicious bots (Ferrara, 2022; Yang and Menczer, 2023).
+In several other cases, scientific findings about bot activity were later found to be consistent with independent platform removals of malicious accounts (Nizzoli et al., 2020; Tardelli et al., 2022), confirming the accuracy of the scientific findings.
+These cases represent just some of the success stories of social bot research.
+Therefore, even if no universal bot detector exists, and despite the many caveats to consider in bot detection, being able to detect some malicious bots puts us in a more advantageous position than being able to detect none.
 Second, the methodological rigor and practical usefulness of some social bots studies are evidenced by corrective actions taken by government agencies based on academic research.
@@
 12, 2014) https://www.sec.gov/resources-for-investors/investor-alerts-bulletins/ia_socialmediafraud following the findings of earlier studies (Hwang et al., 2012).
-Subsequent research confirmed these suspicions in multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020a; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019).
+Subsequent research confirmed these suspicions on multiple occasions (Cresci et al., 2019a; Nizzoli et al., 2020), contributing to calls for greater government oversight of social bots (Gorwa and Guilbeault, 2020; Yan et al., 2023) and discussions on regulating bot freedom of expression (Lamo and Calo, 2019).
 Both the 2018 and 2022 EU Code of Practice on Disinformation, as well as investigations by the U.S.
-Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee have also been informed by research on social bots.
+Congress and the UK Parliament’s Digital, Culture, Media and Sport Committee, have also been informed by research on social bots.
 Furthermore, the concern over bot infiltration was highlighted when Elon Musk initially decided to abandon his purchase of Twitter, citing misleading information about the platform’s financial health and bot prevalence (Varol, 2023)—a claim also supported by Peiter Zatko, Twitter’s former head of security, who exposed serious deficiencies in the company’s security.
-[^8] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/ Table 2 provides some examples of contributions in the study and mitigation of social bots that have withstood the test of time.
-These examples underscore the enduring impact and significance of rigorous social bots research in shaping policy and industry practices.
+[^7] What the Twitter Whistleblower Disclosure Means for Elon Musk: https://time.com/6207992/twitter-bots-elon-musk/
 Third, the benefits of social bot research extend beyond the detection of malicious bots.
@@
 In addition, social bot research has contributed to the early development of other neighboring fields.
-The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience with related phenomena, thereby offering crucial insights when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years.
+The early research on social bots, which dates back to 2010 (Cresci, 2020), provided an important foundation, allowing the field to draw upon several years of accrued experience when widespread concerns regarding misinformation, state-backed trolls, and coordinated inauthentic behavior surged in 2016 and subsequent years.
 In other words, early results on detecting and characterizing social bots informed strategies for detecting and mitigating other related forms of online manipulation.
 In light of these considerations, research on social bots—imperfect as it is—seems far from useless.
-In a sense, research on social bots is no exception to a core framework for the advancement of science: The accumulation of knowledge as a result of research efforts, other than solving only the specific problems the research is addressed to, also fertilizes the scientific ecosystem in many ways (e.g., forming generations of scientists) and provides tools, data, methodology, and insights for scientific advancements in close (and sometimes not so close) research fields.
+In fact, social bots research perfectly exemplifies the process leading the general advancement of science: the accumulation of knowledge resulting from certain research efforts, other than solving local problems, also fertilizes the broader scientific ecosystem, providing data, methodologies, tools, and insights for further scientific advancements in close—and sometimes not so close—fields.
 4 Challenges in the post-API era
@@
 In the API era, most social bot studies have focused on Twitter/X due to its free API and the ease of large-scale data acquisition.
-This hyperfocus on Twitter/X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores platforms beyond Twitter/X.
-However, this data policy ceased in 2023 when Twitter/X terminated free data access for researchers.
+This hyperfocus on X has led to the neglect of other platforms, creating a bias in social bots research that largely ignores other platforms.
+However, this data policy ceased in 2023 when X terminated free data access for researchers.
 [^8] https://www.theverge.com/2023/5/31/23739084/twitter-elon-musk-api-policy-chilling-academic-research Similarly, Reddit restricted free data access, and Meta announced the discontinuation of CrowdTangle, a crucial tool for researchers studying Facebook and Instagram.
 [^9] https://www.niemanlab.org/2024/03/a-window-into-facebook-closes-as-meta-sets-a-date-to-shut-down-crowdtangle
-The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence in real time.
+The lack of access to fresh social media data severely hampers researchers’ ability to monitor bot activities and assess their influence.
 It significantly impedes the collection of new bot samples necessary for studying their characteristics and training novel machine learning classifiers.
@@
 Their opaque and stringent application review processes also leave their efficacy in question (Jaursch et al., 2024).
-Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduce new dynamics.
+Besides traditional social media, the rise of decentralized platforms such as Mastodon and BlueSky introduces new dynamics.
 Their openness facilitates data access for researchers but also exposes them to exploitation by malicious actors.
 The decentralized nature of these platforms complicates efforts to combat threats such as malicious social bots, presenting novel challenges to both users and the research community.
-Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth further contributes to hindering bot detection research on these new platforms.
+Moreover, the limited number of participants, which has not yet reached a critical mass, and the difficulty in creating a reliable ground truth, further contribute to hindering bot detection research on these platforms.
 Overall, the new data accessibility landscape and the change in the underlying technologies, represent an opportunity and a call for further research in social bot detection.
@@
 The way in which new findings are presented in this and in neighboring fields can influence not only the next iterations of research, but also industry practices, policymaking, and public opinion.
-To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that haunt the social bot literature and that currently fuel ambiguities, generalized misunderstandings, and friction among scholars.
+To this end, it is paramount to avoid repeating the mistakes and propagating the common misconceptions that currently haunt the social bot literature and that fuel ambiguities, generalized misunderstandings, and friction among scholars.
 The everyday challenges that we face as researchers in the broad area of misinformation are an accurate reflection of those faced by our society.
 As authors, reviewers, and readers of new research in this field, we have the moral obligation of refraining from falling for the same biases, and exacerbating the same issues, that we frequently encounter in our analyses.
-Misinformation often surfaces as «an accurate fact set in misleading context» (Starbird, 2019).
+Misinformation often surfaces as «an accurate fact set in a misleading context» (Starbird, 2019).
 Perpetuating the methodological and conceptual issues addressed in this paper contributes to creating flawed or unreliable research.
@@
 6 Acknowledgments
-This work was partially supported by project SERICS (PE00000014) under the NRRP MUR program funded by the EU – NGEU; by the European Union – Horizon 2020 Program under the scheme “INFRAIA-01-2018-2019 – Integrating Activities for Advanced Communities”, Grant Agreement n.
-871042, “SoBigData++: European Integrated Infrastructure for Social Mining and Big Data Analytics”; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.
+This work was partially supported by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006) and for the ERC project DEDUCE under grant #101113826.; by project DESIRE 2.0 (DissEmination of ScIentific REsults) funded by IIT-CNR; by the Volkswagen Foundation as part of the Bots Building Bridges project, under the “Artificial Intelligence and the Society of the Future” initiative; by DARPA (awards W911NF-17-C-0094 and HR001121C0169); and by the Knight Foundation.